### v1

# MASQUE: A Text-Guided Diffusion-Based Framework for Localized and Customized Adversarial Makeup

Abstract
As facial recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights. To counteract, various anti-facial recognition techniques have been proposed for privacy protection by adversarially perturbing face images, among which generative makeup-based approaches are the most popular. However, these methods, designed primarily to impersonate specific target identities, can only achieve weak dodging success rates while increasing the risk of targeted abuse. In addition, they often introduce global visual artifacts or a lack of adaptability to accommodate diverse makeup prompts, compromising user satisfaction. To address the above limitations, we develop MASQUE, a novel diffusion-based framework that generates localized adversarial makeups guided by user-defined text prompts. Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity. Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity and stronger adaptability to various text makeup prompts.
1 Introduction
Facial recognition (FR) systems [22] have been adopted in a wide range of security, biometrics, and commercial applications. However, their unregulated deployment poses serious privacy risks, allowing for unauthorized surveillance and malicious tracking. To address these concerns, anti-facial recognition (AFR) technologies [34] have emerged to protect user privacy from unauthorized FR systems. AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online. Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance. Traditional adversarial approaches, such as noise-based methods [13, 37, 41], obscure facial features with norm-bounded global perturbations, while patch-based techniques [36, 15] optimize adversarial patterns in localized image regions. However, these methods often introduce noticeable visual artifacts, compromising the usability of the produced images.
To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial nets or diffusion models for AFR. Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features. A notable direction is makeup-based AFR [38, 12, 26, 30, 7], which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Appendix A for detailed discussions of related work).
While makeup-based AFR methods offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts. In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse. When considering the more privacy-critical dodging scenarios, their performance in protection success rates also drops significantly.
Contribution. We propose MASQUE, a novel diffusion-based image editing method for localized adversarial makeup generation with customized text guidance. In particular, MASQUE is designed to meet the following desiderata simultaneously:
  • Inoffensive Identity Protection: ensure high protection success rates in dodging scenarios without requiring images of a specific target identity.
  • Localized Modification: restrict adversarial perturbations to designated areas, preserving all other facial details.
  • User Control: demonstrate strong prompt-following ability and adaptability to diverse text makeup prompts.
2 Problem Formulation
In this section, we specify the threat model for unauthorized facial recognition (Section 2.1), introduce the background on anti-facial recognition (Section 2.2), and explain why dodging is more suitable than impersonation (Section 2.3).
2.1 Threat Model
We consider the adversarial setup, where an attacker adopts unauthorized FR models to identify benign users from their publicly shared facial images. This enables the extraction of sensitive information, posing significant privacy risks. Since many well-trained FR models are readily available—either as open-source implementations or through commercial APIs—attackers can easily obtain automated FR tools to achieve this malicious objective. Formally, let $\mathcal{X}$ be the input facial image space and $\mathcal{Y}$ be the set of possible identities. Given a collection of online-scraped face images $\mathcal{D}$, the adversary aims to correctly identify victim users from their facial images according to the following objective:
[…]
where $x$ stands for a face image, $y$ is the corresponding ground-truth identity, and $\mathrm{FR}:\mathcal{X}\rightarrow\mathcal{Y}$ denotes a FR model.
A typical FR model consists of three key components: a feature extractor, a gallery database, and a query matching step (see [34, 7] for detailed descriptions). Aligned with prior work [26, 30], we assume that FR models employed by the attacker can achieve high accuracy on clean images. Since FR systems can vary in design, the attacker may utilize different feature extractors and query-matching schemes. To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs. Additionally, we consider face verification as a specialized FR variant, where the gallery database contains only a single user.
2.2 Anti-Facial Recognition
The primary goal of AFR is to evade the identification of unauthorized facial recognition models used by adversaries described in Section 2.1, improving the protection of benign users’ privacy. In this work, we focus on the most popular adversarial-based AFR techniques, which craft imperceptible or naturalistic perturbations (e.g., adversarial makeup) to user’s images to fool FR systems. Specifically, the objective of adversarial-based AFR can be cast into a constrained optimization problem as follows:
$\max\>\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}(x)\big{)}\neq y\big{\}}$ (2)
[…]
where $\mathrm{AFR}:\mathcal{X}\rightarrow\mathcal{X}$ denotes the AFR perturbation function, $\Delta$ is a similarity metric that measures the visual distortion of the AFR-perturbed image $\mathrm{AFR}(x)$ with reference to the original image $x$, and $\gamma>0$ is a threshold parameter reflecting the distortion upper bound the user can tolerate.
Dodging Success. We consider benign users who seek privacy protection through adversarial-based AFR before sharing their facial images online. A desirable AFR technique should achieve a high dodging success rate (DSR) against unauthorized facial recognition, as defined by the optimization objective in Equation LABEL:eq:adversarial_AFR. For clarity, we refer to these users as defenders, as they apply AFR techniques to modify their facial images with adversarial perturbations. Since the defender typically lacks precise knowledge of the FR model employed by the attacker, we consider the black-box scenario when evaluating the DSR of each AFR method.
Visual Quality. From the user’s perspective, images perturbed by AFR should appear natural and closely resemble the original, as captured by the optimization constraint in Equation LABEL:eq:adversarial_AFR. If $\mathrm{AFR}(x)$ has an unsatisfactory image quality, users may be reluctant to share the protected face image online, even if a high DSR is achieved. Given this expectation of high visual quality, earlier privacy-preserving techniques such as face obfuscation or anonymization [20, 29] are not suitable. In comparison, makeup-based AFR methods [26, 30] stand out, as they introduce natural adversarial perturbations with minimal visual distortion. In this work, we further improve the performance of generative makeup-based AFR, enhancing both DSR and visual quality.
Figure 1: Each row presents the reference image with a text makeup prompt, the original image, followed by the generations of each AFR method. Methods marked with $\dagger$ require only a text prompt for makeup application and do not rely on a reference image. The yellow text indicates the cosine similarity score. Below the images, we compare ISR and DSR averaged across four FR models and three makeup styles.
2.3 Why Dodging instead of Impersonation?
In contrast to the untargeted dodging objective, several state-of-the-art studies [37, 12, 26, 30] have focused on improving AFR performance in targeted impersonation scenarios. Impersonation-oriented methods aim to adversarially modify a user’s facial images to mislead a FR model into recognizing them as a specific target identity. Let $\mathrm{AFR}_{\mathrm{I}}:\mathcal{X}\rightarrow\mathcal{X}$ be the perturbation function for an impersonation-oriented method. In existing literature, the impersonation success rate (ISR) is typically measured by the following metric:
$\text{ISR}:=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}_{\mathrm{I}}(x)\big{)}=y_{t}\},$ (3)
where $y_{t}$ denotes the target identity satisfying $y_{t}\neq y$. Below, we further detail why impersonation is not suitable for AFR.
Misaligned Goal. From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics, ISR and DSR. As shown by [42], high ISR does not guarantee high DSR. If the target identity $y_{t}$ is not included in the attacker’s model (e.g., using facial verification with only a single image of the victim identity $y$), optimizing for higher ISR offers little benefit due to misaligned objectives. As we will illustrate in Section 3.1, prior impersonation-oriented AFR methods can only achieve weak privacy protection when evaluated under dodging scenarios. Therefore, we argue that effective privacy protection should prioritize untargeted dodging success.
Targeted Misuse. Impersonation increases the risk of targeted misuse, raising serious ethical and fairness concerns. Unlike dodging, impersonation-based approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse. This ability can be exploited for malicious purposes, such as falsely associating one’s actions with another individual. For instance, if a user $y$ intentionally applies an AFR tool to impersonate a target $y_{t}$ and engages in illegal activity, FR systems may misidentify $y_{t}$ as the suspect. This deliberate identity manipulation poses significant ethical and legal risks [34].
By contrast, dodging-oriented AFR serves as a better solution for facial privacy protection. Instead of redirecting recognition to a specifically chosen identity, dodging causes misclassification in an uncontrolled manner, typically as an unknown or incorrect identity. This lack of control makes dodging significantly less susceptible to intentional misuse, as users cannot deliberately impersonate or frame a different party. While dodging still prevents correct identification, it does so without assigning a specific false identity, reducing the risk of targeted deception or malicious intent.
Dodging External ID Guidance Local Prompt-Following
TIP-IM [37] n/a n/a
AMT-GAN [12] image low
C2P [26] text medium
DiffAM [30] image medium
MASQUE (ours) text high
Table 1: Comparisons of SOTA AFR methods for key features in terms of both privacy protection and visual quality.
3 Challenges of Existing AFR Methods
In this section, we present preliminary results to highlight key limitations of previous work, further motivating the design of our approach, which will be detailed in Section 4.
Preliminary Results. Figure 1 shows qualitative comparisons of three makeup styles, evaluating several prior AFR methods, including a noise-based method, TIP-IM [37], and three state-of-the-art makeup-based approaches [12, 26, 30]. For each AFR method, we report the metrics across different black-box facial verification models averaged across $100$ randomly selected images from CelebA-HQ [14]. To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset [16] or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 for more details). Table 1 summarizes the key feature comparisons of state-of-the-art AFR methods, including ours.
3.1 Weak Protection under Dodging
As discussed in Section 2.3, an ideal AFR should account for dodging scenarios, to ensure privacy without dependence on a target identity. However, despite its ethical advantages, dodging remains an underexplored and underdeveloped area in makeup-based AFR. Existing methods that perform well under impersonation settings often exhibit a significant drop in DSR, as shown in Figure 1. This highlights a fundamental gap: current techniques are not optimized for dodging, leaving privacy protection less reliable when a user seeks to avoid recognition rather than impersonate another identity. In addition, while some methods like [26, 7] introduce a form of dodging, they still rely on a target identity. This reliance contradicts the very goal of dodging-based AFR, which should function without any identity substitution. Thus, a key challenge in AFR research is developing a robust dodging-based approach that provides strong protection while avoiding the ethical risks associated with impersonation.
3.2 Non-Localized Face Modification
Noise-based approaches like TIP-IM generate global perturbations, which fail to localize the editings. Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches, including AdvMakeup, AMT-GAN, and C2P, introduce global artifacts that alter non-facial regions. This lack of precise localization is particularly problematic when users seek to protect only their facial identity while preserving the background. As shown in Figure 1, this issue is especially pronounced in complex backgrounds, where perturbations should remain confined to facial regions but instead extend beyond. In contrast, the diffusion-based method DiffAM enables better localization, effectively applying makeup perturbations from a reference image. However, its reliance on reference images limits flexibility, particularly when suitable references are unavailable. This highlights the need for methods that achieve both precise localization and greater flexibility.
3.3 Limited Prompt-Following Ability
Makeup-based AFR methods aim to modify facial appearance to prevent accurate identification by FR systems. However, existing methods offer limited user control, making customization and usability a key challenge. Broadly, makeup-based AFR methods fall into two categories:
  • Image-reference-based AFR transfers makeup styles from a reference image. While they aim for accurate replication, they struggle with reliability and flexibility. As shown in Figure 1, early GAN-based methods like AMT-GAN fail to transfer makeup consistently, while DiffAM, despite improvements, remains restricted. It only applies makeup to three fixed regions—skin, mouth, and eyes—preventing transfer to other areas like colored eyebrows. Additionally, these methods require separate fine-tuning for each reference, limiting efficiency.
  • Text-prompt-based AFR allows users to describe makeup styles in natural language, eliminating the need for reference images or separate fine-tuning while enhancing privacy. However, this method lacks fine-grained control, as C2P relies solely on CLIP directional loss, often leading to unintended modifications beyond the intended areas.
4 Our Methodology: MASQUE
Figure 2: The pipeline of MASQUE involves: (1) fusing the editing and reconstruction prompts to produce an updated cross-attention map for diffusion, (2) creating a mask $\mathcal{M}$ to define a target region and optimize an edit loss to maximize makeup-related attention in $\mathcal{M}$, and (3) using pairwise adversarial guidance with same-identity image to enhance identity confusion without external targets.
To address the key challenges discussed in Section 3, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced. Figure 2 illustrates the pipeline of MASQUE with its pseudocode in Appendix D. The makeup is guided by a user-defined text prompt $p^{*}$, refined via cross-attention fusion with a mask. Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR systems without compromising visual fidelity.
Our method leverages the state-of-the-art Stable Diffusion [24] adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent $z_{T}$ conditioned on a text embedding $C$. To be more specific, Stable Diffusion is trained to predict the added noise $\epsilon$ via:
$\min_{\theta}\ \mathbb{E}_{{z}_{0},\epsilon\sim\mathcal{N}(0,I),t\sim\mathrm{Unif}(1,T)}\bigg{[}\big{\|}\epsilon-\epsilon_{\theta}(z_{t},t,C)\big{\|}_{2}^{2}\bigg{]}.$
Applying Makeup with Cross Attention. Before applying edits, we first obtain a faithful latent representation of the original image $x$ using null-text inversion [19], which mitigates reconstruction errors common in direct DDIM inversion [28]. Conditioning on an empty prompt to align the forward and reverse diffusion trajectories, ensuring near-perfect reconstruction of the image’s structure and identity.
With this accurate latent representation, we introduce makeup attributes via the text prompt $p^{*}$ by manipulating the cross-attention (CA) layers of the diffusion model [24], which control how spatial features correspond to semantic tokens. At each diffusion step $\tau$, we extract attention maps $A_{\tau}$ (reconstruction) and $A_{\tau}^{*}$ (editing) and blend them: preserving CA values from $A_{\tau}$ for shared tokens to maintain structure, while incorporating values from $A_{\tau}^{*}$ for makeup-specific tokens in $p^{*}$:
$\big{(}\mathrm{Update}(A_{\tau},A_{\tau}^{*})\big{)}_{i,j}:=\begin{cases}(A_{\tau})_{i,j},&\text{if $j$ is in both $p$ and $p^{*}$},\\[6.0pt] (A_{\tau}^{*})_{i,j},&\text{if $j$ is unique to $p^{*}$},\end{cases}$
where $p$ denotes the original text prompt. The result is $\hat{A}_{\tau}$, a set of mixed CA maps that preserve the original facial layout while steadily introducing adversarial makeup features [9].
Enhancing Semantic Edits and Locality. To ensure precise localization, we generate a mask $\mathcal{M}$ that defines the region for modification. To achieve this, we embed the prompt $p^{*}$ using a Sentence Transformer [23] model and compare it to embeddings of predefined facial regions. The closest match determines the relevant area for the edit. For instance, if the prompt specifies “a face with red lipstick”, the model identifies lips as the target and generates a lip-area mask. This ensures that our adversarial makeup perturbations remain focused on the correct facial region and do not unintentionally alter other parts of the face.
Once the target region $\mathcal{M}$ is determined, we enhance the influence of makeup-related tokens by maximizing their attention within $\mathcal{M}$. This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality [18]. Specifically, we optimize:
$\mathcal{L}_{\mathrm{edit}}=\left(1-\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\frac{(A_{\tau}^{*})_{i,\mathrm{new}}}{(A_{\tau}^{*})_{i,\mathrm{new}}+\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}}\right)^{2},$
where $(A_{\tau}^{*})_{i,\mathrm{new}}$ is the attention weights assigned to the new makeup tokens at spatial index $i$, while $\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}$ is the total attention weight of tokens that appear in both the original and makeup prompts. By emphasizing makeup-specific attention in the masked region, this loss term ensures that modifications in $x_{\mathrm{p}}$ occur meaningfully in the intended region while preventing edits from spreading elsewhere.
To further enforce spatial precision, we process the latent through two distinct branches: an edit branch for makeup application and a reconstruction branch to preserve original features. During the backward steps, we explicitly constrain perturbations within the designated area by imposing the following conditions:
$z_{\text{edit}}=\mathcal{M}\cdot z_{\text{edit}}+(1-\mathcal{M})\cdot z_{\text{rec}},$
where $z_{\text{edit}}$ represents the latent from the edit branch, and $z_{\text{rec}}$ denotes the latent from the reconstruction branch.
While makeup edits ensure semantic plausibility, the core adversarial objective is to disrupt FR models. Here lies our novel contribution: a pairwise adversarial guidance that uses a guide image of the same identity to achieve robust identity confusion without relying on an external target.
Method CelebA-HQ VGG-Face2-HQ Avg.
IR152 IRSE50 FaceNet MobileFace IR152 IRSE50 FaceNet MobileFace
Identification Clean $10.00$ $13.00$ $5.00$ $40.00$ $13.00$ $18.00$ $18.00$ $25.00$ $17.75$
TIP-IM $62.00$ $86.00$ $65.00$ $74.00$ $57.00$ $73.00$ $52.00$ $62.00$ $66.38$
AMT-GAN $61.22$ $48.33$ $50.67$ $57.33$ $16.00$ $24.00$ $22.33$ $32.33$ $39.03$
DiffAM $54.00$ $57.67$ $59.00$ $74.33$ $49.67$ $51.67$ $54.00$ $70.67$ $58.88$
C2P (I) $30.67$ $38.33$ $22.00$ $56.67$ $18.00$ $20.00$ $21.67$ $32.00$ $29.92$
C2P (D) $74.67$ $76.33$ $56.33$ $77.33$ $18.33$ $19.67$ $20.33$ $30.33$ $46.67$
MASQUE ($G=0$) $92.33$ $95.67$ $61.33$ $87.00$ $69.67$ $76.33$ $45.33$ $81.67$ $76.17$
MASQUE ($G=1$) $\underline{98.00}$ $\underline{98.33}$ $\underline{77.67}$ $\underline{94.00}$ $\underline{84.33}$ $\underline{89.00}$ $\underline{61.00}$ $\underline{91.00}$ $\underline{85.50}$
Verification Clean $5.00$ $5.00$ $4.00$ $10.00$ $13.00$ $17.00$ $15.00$ $30.00$ $12.38$
TIP-IM $44.00$ $60.00$ $51.00$ $40.00$ $46.00$ $51.00$ $59.00$ $46.00$ $49.63$
AMT-GAN $40.00$ $32.00$ $51.67$ $25.67$ $14.00$ $26.33$ $26.67$ $40.67$ $32.13$
DiffAM $31.00$ $26.00$ $54.00$ $43.33$ $46.33$ $50.67$ $56.00$ $82.67$ $48.75$
C2P (I) $12.33$ $11.33$ $14.33$ $21.33$ $16.33$ $18.67$ $19.33$ $37.33$ $18.88$
C2P (D) $52.33$ $46.67$ $40.33$ $41.67$ $17.00$ $19.67$ $18.33$ $38.00$ $34.25$
MASQUE ($G=0$) $89.33$ $90.67$ $57.67$ $76.33$ $62.33$ $71.33$ $42.33$ $86.00$ $72.00$
MASQUE ($G=1$) $\underline{96.00}$ $\underline{96.67}$ $\underline{75.67}$ $\underline{79.33}$ $\underline{82.33}$ $\underline{91.67}$ $\underline{63.33}$ $\underline{94.33}$ $\underline{84.92}$
Table 2: Dodging success rate (%) under black-box facial identification and verification settings. For each column, the other three FR systems are used as surrogates to generate the protected faces, and the best result is underlined. Here, $G$ denotes the number of guide images.
Pariwise Adversarial Guidance. Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios. By contrast, our approach leverages a pair $(x,\tilde{x})$ of face images from the same individual, where $\tilde{x}$ serves as the guide image. This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss. Note that in the standard diffusion denoising process, the objective is to generate images similar to the original. If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness (see Tables 2-3 in Appendix B for supporting evidence and comparison results with self-augmented guide images).
Adversarial and Image Quality Trade-off. We introduce adversarial perturbations during the later stage of the diffusion process, ensuring coarse structure remains intact while subtly altering identity-specific features. To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure that captures adversarial effectiveness:
$\mathcal{L}_{\mathrm{adv}}=\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}(x_{\mathrm{p}},\tilde{x})+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(x_{\mathrm{p}},x),$
where $x_{\mathrm{p}}$ denotes the protected image. $\mathrm{CoSi}$ ensures that the adversarial perturbations sufficiently diverge from recognizable identity features, while $\mathrm{LPIPS}$ [39] maintains perceptual and structural fidelity, respectively. The parameters $\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$ allow fine-tuning of this trade-off. We conduct ablation studies to understand the sensitivity of these hyperparameters and how to set them in MASQUE for better trade-offs (see Section 5.3 and Appendix B).
5 Experiments
In this section, we conduct comprehensive experiments to study the effectiveness of MASQUE with comparisons to prior AFR methods in both privacy protection and visual aspects. Appendices B and C present additional experimental results and visualizations for more general makeup prompts.
5.1 Experimental Settings
Dataset. To reflect the high-quality nature of online facial images, we conduct experiments using $300$ images from each of the CelebA-HQ [14] and VGG-Face2-HQ [3] datasets, all at a resolution of $1024\times 1024$. We randomly sample $100$ identities from per dataset, each with three images: a probe image to be protected, a reference image for guidance, and a the gallery image, which is assumed to be stored in the facial recognition system for comparison.
Configuration. We compare our methods against the following baselines: TIP-IM [37], the SOTA noised-based AFR, and other recent generative-based AFR methods, including AMT-GAN [12], Clip2Protect [26], and DiffAM [30]. For Clip2Protect, we denote it as C2P for simplicity, where both the impersonating version (denoted as C2P (I)) and the dodging version (i.e., C2P (D)) are assessed. We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 [6], IRSE50 [11], FaceNet [25], and MobileFace [2]. We also evaluate AFR methods against commercial FR APIs: Face++  [^1] https://www.faceplusplus.com/face-comparing/ and Luxand  [^2] https://luxand.cloud/face-api.
(a) Red Lipstick
(b) Blue Eyebrow
(c) Pink Eyeshadow
Figure 3: Comparison of concentration and hue accuracy across different methods for three makeup tasks: red lipstick, blue eyebrow, and pink eyeshadow. Concentration values are normalized relative to MASQUE.
For reference-based methods, we selected images from the makeup dataset [16] used during their pre-training, that best matched the given prompt. For the blue eyebrow style—introduced to assess performance on an uncommon makeup—finding a suitable reference was challenging, underscoring the limitations of reference-based methods for rare styles. To address this, we inpainted blue eyebrows on a non-makeup image from the dataset using our framework, excluding pairwise adversarial guidance, demonstrating its capacity for targeted edits without an external reference.
Evaluation Metric. We employ dodging success rate (DSR) as the primary evaluation metric for AFR. DSR is computed using a thresholding strategy for face verification and a closed-set strategy for face identification. For verification, specifically, DSR is defined as:
$\mathrm{DSR}=\frac{1}{N}\sum_{x}\mathbb{I}\big{(}\cos(\mathrm{FR}(x_{\mathrm{p}}),\mathrm{FR}(x_{\mathrm{g}}))>\tau\big{)}\times 100\%,$
where $\mathbb{I}$ is the indicator function, $N$ is the number of face images and $\mathrm{FR}$ represents the target face recognition model. $x_{\mathrm{p}}$ denotes the protected image, while $x_{\mathrm{g}}$ stands for the gallery image of the corresponding identity. The similarity threshold $\tau$ is set at a $0.01$ False Acceptance Rate (FAR) for each victim model. For face identification, we use the Rank-1 Accuracy, which measures whether the top-1 candidate list excludes the original identity of $x$.
To assess the visual quality, we use LPIPS, PSNR, and SSIM [33]. A lower LPIPS score reflects higher perceptual similarity, while higher PSNR and SSIM values indicate better pixel-wise and structural fidelity to the original image.
Implementation Detail. MASQUE builds on the pre-trained Stable Diffusion v1.4 model, using DDIM denoising over $T=50$ steps with a fixed guidance scale of $7.5$. During the backward diffusion process, CA injection occurs in $[T,T-\tau_{\mathrm{attn}}]$ ($\tau_{\mathrm{attn}}=40$), localize optimization in $[T,T-\tau_{\mathrm{edit}}]$ ($\tau_{\mathrm{edit}}=5$), and adversarial guidance in $[T-\tau_{\mathrm{adv}},0]$ ($\tau_{\mathrm{adv}}=45$). We set $\lambda_{\mathrm{CoSi}}=0.1$, $\lambda_{\mathrm{LPIPS}}=1$, and cap optimization to $\mathrm{max}\_\mathrm{adv}=15$ iterations.
5.2 Main Comparison Results
Against Black-Box FR Models. Table 2 presents the DSR for both face verification and Rank-1 identification in black-box scenarios, evaluated using four widely used pre-trained FR feature extractors [6, 11, 25, 2]. For each target model, the remaining three serve as surrogate models to simulate the black-box scenario, with results averaged across three makeup styles. Our method significantly outperforms all baselines, achieving an average DSR of $85.50\%$ for identification and $84.92\%$ for verification. In addition, while MASQUE without a guide image achieves a reasonable DSR, it still falls short compared to its guided counterpart. Furthermore, image quality degrades significantly without a guide image (see supporting evidence in Appendix B.2.).
(a) Face++
(b) Luxand
Figure 4: Performance comparisons against commercial FR APIs. Confidence scores (lower is better) from Face++ and Luxand show that MASQUE provides stronger privacy protection than state-of-the-art noise-based and makeup-based AFR approaches.
Real-world Performance. We also test our method against the commercial APIs, Face++ and Luxand, in verification mode, which assigns similarity scores from 00 to $100$. Figure 4 shows the results. As a proprietary model with unknown training data and parameters, it serves as a realistic testbed for evaluating the effectiveness of AFR methods. Our approach achieves the lowest similarity score across both APIs, demonstrating effectiveness in both open-source and closed-source settings, reinforcing its real-world applicability.
TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE
LPIPS (↓) $0.311$ $0.342$ $0.460$ $0.473$ $0.399$ $0.294$
PSNR (↑) $32.158$ $19.512$ $18.922$ $17.990$ $18.314$ $25.815$
SSIM (↑) $0.928$ $0.613$ $0.583$ $0.563$ $0.769$ $0.856$
Table 3: Quantitative comparisons of the visual quality across different AFR methods. Lower values are better for LPIPS, while higher values are better for PSNR and SSIM.
Visual Quality. Our method achieves superior image quality across multiple evaluation metrics, as summarized in Table 3 and illustrated in Figure 1. While TIP-IM attains the highest PSNR and SSIM due to its small perturbation budget, these pixel-level metrics often fail to reflect perceptual quality. In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance. We also confirm the advantages of MASQUE over existing AFR methods in achieving precise, localized makeup transfer (see additional experiments in Appendix B.1).
Prompt Adherence. We assess prompt adherence in terms of spatial precision and color accuracy. For spatial adherence, we use Concentration, which quantifies the proportion of modifications within a binary mask of the target area. Higher scores indicate better localization, while lower scores suggest spillover. For color precision, we evaluate Hue Accuracy by extracting modified pixels within the mask, converting them to HSV space, and measuring how many fall within the expected hue range. Figure 3 shows that MASQUE consistently achieves the highest concentration scores and competitive hue accuracy across prompts, ensuring precise localization with accurate color application. In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility. These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
(a) $\tau_{\mathrm{adv}}$
(b) $\mathrm{max}\_\mathrm{adv}$
Identification DSRVerification DSRLPIPS
Figure 5: Ablation studies on hyperparameters $\tau_{\mathrm{adv}}$ and $\mathrm{max}\_\mathrm{adv}$, highlighting the trade-off between DSR and image quality.
5.3 Ablation Studies
Adversarial Strength vs. Image Quality. Increasing $\tau_{adv}$ enhances the adversarial effect, improving DSR but degrading image quality, as indicated by higher LPIPS values in Figure 5. Similarly, increasing $\mathrm{max}\_\mathrm{adv}$ amplifies protection but further compromises image quality.
Customized Makeup. Figure 6 demonstrates the controllability of MASQUE across the following three distinct aspects. Additional visualizations are available in the Appendix C.
Sequential Editing. MASQUE enables progressive makeup application, allowing users to apply makeup sequentially to multiple facial areas. Each additional layer of makeup enhances the protection effect, as reflected in the decreasing similarity scores shown in Figure 6.
Intensity Control. The intensity of the makeup can be adjusted via text prompts. A simple modifier like “vivid” alters the cross-attention, resulting in color changes. Intensity variation has minimal effect on adversarial effectiveness.
Style Control. MASQUE supports full-face transformations. Prompts like “tanned skin” can blend naturally, while dramatic styles like “clown makeup” are also well-adapted. Greater makeup coverage further strengthens protection.
Figure 6: Examples showing the controllability of MASQUE through text prompts. The number at the bottom right indicates the similarity score between the image and the corresponding gallery image.
6 Further Discussions
Guide Images. While MASQUE with guide images enhance DSR and image quality, they may not always be available. As an alternative, we test self-augmented inputs as guide images via flipping and translation. However, preliminary results in Appendix B show limited improvements, indicating the need for more effective strategies to substitute the guide image.
Computational Efficiency. MASQUE enables adversarial makeup applications without pretraining. However, due to the iterative nature of diffusion models, inference speed remains a computational bottleneck. Future work can explore acceleration schemes to improve the efficiency of MASQUE.
Adaptive Evaluations. Real-world AFR threats evolve, with attackers refining strategies over time [7]. Considering adaptive scenarios, such as post-processing transformations (e.g., compression, smoothing, and makeup removal) and adversarial fine-tuning, could help improve robustness and provide a more comprehensive evaluation of AFR methods.
7 Conclusion
We introduced MASQUE, a diffusion-based AFR framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks. Unlike prior AFR methods, MASQUE achieves high DSR while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse. Our experiments demonstrated the strong performance of MASQUE, where we highlight the important role of pairwise adversarial guidance. Future work may focus on real-time optimization and robustness against evolving FR models to further strengthen facial privacy protection.
Availability
The code implementation of MASQUE and our experiments are open-sourced and available at https://github.com/TrustMLRG/MASQUE.
Appendix A Related Work
A.1 Anti-Facial Recognition
Earlier works on AFR proposed to use obfuscation techniques to obscure the facial identity features [20] or craft $\ell_{p}$-norm bounded perturbations to fool FR models [37]. While effective, these methods often compromise image quality, limiting their practicality in real-world scenarios. Poisoning-based methods [4, 27] introduced a new approach by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models. These techniques excel at disrupting model training or inference without visibly altering the image, but their reliance on model-specific perturbations limits their generalizability across diverse architectures. Unrestricted adversarial examples, which aim to create natural-looking modifications, marked a significant advancement by integrating adversarial signals into realistic alterations without relying on specific constraints. These strategies leverage the flexibility of generative models to balance visual coherence and adversarial effectiveness.
Adversarial Makeup. Adversarial makeup [38, 12, 26, 30] has emerged as a practical solution to realize the goal of facial privacy protection. These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while largely preserving the aesthetic appeal. For example, [12] proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the generative adversarial network (GAN) framework [8]. Recent advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect [26] and DiffAM [30], which adopt text-guided StyleGAN model and diffusion-based framework, respectively, to enable seamless adversarial face modifications with much improved visual quality. In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for privacy protection in dodging scenarios.
A.2 Diffusion-Based Image Editing
Localized Image Editing. Localized editing methods extend diffusion models for targeted modifications. For example, mask-based approaches [1, 5] constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency. Attention-based approaches [9, 32] guide edits via attention injection, achieving better global structure preservation but suffering from unintended changes (editing leakage). Our proposed method, MASQUE, utilizes both mask-based and attention-based guidance without any fine-tuning, combining mask precision with attention flexibility, ensuring semantically consistent and region-specific modifications while addressing the limitations of prior approaches.
Text-Guided Diffusion Models. Text-guided diffusion models are a major advancement in generative AI, capable of synthesizing high-quality images from natural language descriptions. Built on denoising diffusion probabilistic models (DDPMs) [10], they iteratively refine random noise into coherent images guided by text prompts. Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation. Mask-based approaches [1, avrahami2022blended, 5] achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions. While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios. On the other hand, mask-free attention-based methods [9, cao2023masactrl, 32] use attention injection mechanisms to guide edits without requiring explicit masks. These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region. In this work, we leverage mask-based cross-attention guidance [18] to achieve localized adversarial perturbations in the form of makeup. This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts while addressing the limitations of structural inconsistencies seen in previous methods.
A.3 Makeup Transfer
Advancements in makeup transfer have moved from GAN-based methods to more advanced diffusion models. Traditional approaches like BeautyGAN [16] and PSGAN [17] use histogram matching, attention mechanisms, and spatial encodings to transfer makeup while preserving structure. CPM [21] expands beyond color transfer using UV mapping, while RamGAN [35] ensure component consistency. However, GAN-based models struggle with extreme styles and rely on imprecise pseudo-paired data, limiting fidelity.
To overcome these challenges, diffusion-based methods like SHMT [31] and Stable-Makeup [40] offer improved realism and robustness. Specifically, SHMT eliminates reliance on pseudo-paired data through a self-supervised “decoupling-and-reconstruction” framework, leveraging a Laplacian pyramid for hierarchical texture transfer. Stable-Makeup employs a diffusion model with a Detail-Preserving encoder and cross-attention layers to ensure precise, structure-preserving makeup application. These approaches set a new standard for high-fidelity transfer. However, our study takes a different approach—rather than transferring makeup from a reference image, we focus on text-guided makeup generation using diffusion models, enabling intuitive and flexible image editing to create customizable makeup based solely on textual descriptions. This also eliminates the need for model fine-tuning on specific makeup images, allowing us to generate the desired makeup directly during inference, resulting in a more efficient and adaptable approach.
Appendix B Additional Experiments
B.1 Localized Editing
To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask, isolating the corresponding areas in both images via element-wise multiplication. The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons. Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences.
Metric Method In-Mask Out-Mask $\textbf{Diff}\ \bm{\Delta}$
DISTS ()($\downarrow$)( ↓ ) TIP-IM $2.013$ $0.094$ $1.919$
AMT-GAN $7.993$ $0.149$ $7.844$
C2P (I) $8.945$ $0.153$ $8.792$
C2P (D) $10.011$ $0.169$ $9.842$
DiffAM $9.645$ $0.141$ $9.504$
MASQUE $12.306$ $0.104$ $12.202$
LPIPS ()($\downarrow$)( ↓ ) TIP-IM $0.322$ $0.305$ $0.017$
AMT-GAN $0.360$ $0.316$ $0.044$
C2P (I) $0.400$ $0.447$ $-0.047$
C2P (D) $0.436$ $0.460$ $-0.025$
DiffAM $0.509$ $0.410$ $0.099$
MASQUE $0.434$ $0.290$ $0.144$
PieAPP ()($\downarrow$)( ↓ ) TIP-IM $3.621$ $0.462$ $3.1591$
AMT-GAN $21.325$ $0.814$ $20.511$
C2P (I) $17.186$ $1.138$ $16.049$
C2P (D) $20.850$ $1.314$ $19.536$
DiffAM $17.523$ $1.312$ $16.211$
MASQUE $21.433$ $0.645$ $20.789$
Table 4: Comparison of perceptual similarity metrics across in-mask and out-mask regions and different AFR methods.
Table 4 compares the effectiveness of localized edits, where higher $\Delta$ values indicate stronger localization of perturbations. TIP-IM [37], as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions. In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
B.2 Pairwise Adversarial Guidance
MASQUE introduces pairwise adversarial guidance to protect identities without external target references, distinguishing it from impersonation-oriented methods. By aligning features between the guide and original images, we identify identity-relevant cues and inject subtle adversarial signals that steer the diffusion process away from a recognizable identity manifold. The guide image serves as a proxy for the gallery image, providing guidance on the direction in which the latent representation should be altered. This eliminates reliance on external identities, mitigating ethical concerns related to impersonation while ensuring identity protection remains non-intrusive.
Leveraging multiple guide images further stabilizes adversarial training, reducing bias toward any single representation and improving robustness. To evaluate its effectiveness, we use FaceNet as the target model and select $100$ identities from CelebA-HQ, varying the number of guide images per identity. These evaluation identities are distinct from the main experiments, ensuring at least $12$ images per identity to accommodate setups with up to $10$ guide images.
Effectiveness is measured using the confidence score, the gap between the similarity score of the original identity’s gallery image and that of the misidentified identity. Table 5 shows that increasing guide images enhances identity obfuscation, leading to higher confidence scores while reducing similarity with the original identity. Moreover, images generated with guide images in the pairwise adversarial guidance step exhibit higher visual quality than those relying solely on adversarial loss on the original image. This confirms that simply maximizing distance from the original image is ineffective, as it disrupts the diffusion process’s goal of preserving natural image characteristics, leading to instability in both image quality and adversarial performance. These results demonstrate that MASQUE’s pairwise adversarial guidance provides a robust, privacy-centric solution, balancing strong identity obfuscation with high visual fidelity.
#Guide Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) Confidence()($\uparrow$)( ↑ ) Similarity()($\downarrow$)( ↓ ) LPIPS()($\downarrow$)( ↓ )
00 $0.66$ $0.60$ $0.114$ $0.355$ $0.303$
$1$ $0.67$ $0.63$ $0.135$ $0.349$ $0.292$
$2$ $0.72$ $0.71$ $0.165$ $0.297$ $0.291$
$5$ $0.74$ $0.74$ $0.177$ $0.283$ $0.291$
$10$ $0.76$ $0.77$ $0.197$ $0.266$ $0.291$
Table 5: Comparison of identification confidence, verification similarity, and image quality metrics across different numbers of guide images employed in MASQUE.
Guide Method Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ )
Pairwise ($G=1$) $0.72$ $0.66$ $25.691$
Self-Guidance ($G=0$) $0.56$ $0.40$ $25.812$
    + Horizontal Flip $0.57$ $0.50$ $25.834$
    + Vertical Flip $0.12$ $0.06$ $26.566$
    + Translation $0.55$ $0.38$ $25.807$
Table 6: Impact of adversarial image guidance in DSR and PSNR. We test pairwise- and self-guidance with different augmentations.
B.3 Self-Augmented Guidance
To address cases where users may lack suitable guide images, we tested using self-augmented input images as guide images. This experiment was conducted on $100$ identities from CelebA-HQ with the prompt “red lipstick”, targeting the FaceNet model. For augmentation, we applied horizontal flipping, vertical flipping, and random translation with a maximum shift of $20$ pixels. Results, as shown in Table 6, compare these approaches against the input image itself as the guide image. Horizontal flipping and random translation perform similarly to using the input image as the guide, whereas vertical flipping shows minimal adversarial effect. This is likely because vertically flipped faces are hard to align with facial images by the recognition model.
B.4 DSR under Rank-5 Accuracy
We also report DSR for the face identification task, using rank-5 accuracy to assess dodging success. The results are presented in Table 7. Notably, MASQUE achieves the highest DSR in both rank-1 and rank-5 settings, demonstrating its effectiveness in misleading identification models.
AMT-GAN DiffAM C2P (I) C2P (D) MASQUE
Rank-1 $54.39$ $61.25$ $36.92$ $71.17$ $92.08$
Rank-5 $28.58$ $45.58$ $18.25$ $48.67$ $81.42$
Table 7: DSR with Rank-1 and Rank-5 metrics across AFR methods
B.5 Mask and Perturbation Localization
Figure 7: Visualizations of MASQUE with and without masking. The first row uses prompt “shimmery gold eyeshadow”, while the second row uses “blue eyeshadow”.
To understand the impact of masking in MASQUE, we conduct an ablation study with and without applying a mask. Figure 7 shows that the absence of a mask causes the model to struggle with spatial control, leading to unconstrained perturbations across the entire image rather than confining makeup to the intended area.
Appendix C Additional Visualizations
Figure 8: Visualizations of MASQUE-generated images with or without using the edit loss.
Additional visual results of image-guided makeup transfer are provided in Figure 8 to accommodate space limitations in the main text. MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics.
To evaluate color variation, we provide examples featuring both a male and a female subject. Additionally, by testing prompts that target different facial regions (eyes, lips, cheeks, skin, and full-face makeup), we demonstrate MASQUE’s ability to generate realistic makeup across various regions. We further provide visualizations without $\mathcal{L}_{\text{edit}}$ optimization to illustrate its impact.
Appendix D Algorithm Pseudocode
Input: $\bm{z}_{\mathrm{edit}},\bm{z}_{\mathrm{rec}}$: Edited and original latent, $\bm{e}_{\mathrm{edit}},\bm{e}_{\mathrm{rec}}$: Edited and original text embeddings, $\mathcal{D}$: Stable Diffusion model, $T$: Number of diffusion steps, $\tau_{\mathrm{attn}},\tau_{\mathrm{edit}},\tau_{\mathrm{adv}}$: Step thresholds, $\mathcal{M}$: Binary region mask, $\mathcal{I}_{\mathrm{target}}$: Target token indices, $d_{k}$: Key/Query dimensionality, $\lambda_{\mathrm{edit}},\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$: Loss weights, $\eta$: Learning rate, $\sigma(t)$: Noise scale function, $\bm{x}$: Input image, $\tilde{\bm{x}}$: Guide image, $max\_adv$: Maximum adversarial iterations
Output: Final protected image $\bm{x}_{\mathrm{p}}$
for $k\leftarrow T$ to $1$ do
      if $k>T-\tau_{\mathrm{attn}}$ then
             //Perform cross-attention edit
             $\mathrm{CA}^{\mathrm{refined}}\leftarrow\{\}$
            for $\ell\in\text{CrossAttnLayers}$ do
                   $Q_{\ell},K_{\ell},V_{\ell}\leftarrow\mathcal{D}.\mathrm{UNet.GetCrossAttentionComponents}(\bm{z}_{\mathrm{edit}},\bm{e}_{\mathrm{edit}},\ell)$
                  $\mathrm{CA}_{\ell}\leftarrow\mathrm{Softmax}\Bigl{(}\frac{Q_{\ell}\,K_{\ell}^{\top}}{\sqrt{d_{k}}}\Bigr{)}$
                  $\mathrm{CA}_{\ell}^{\mathrm{refined}}\leftarrow\mathrm{CA}_{\ell}\cdot V_{\ell}$
                  $\mathrm{CA}^{\mathrm{refined}}\leftarrow\mathrm{CA}^{\mathrm{refined}}\cup\{\mathrm{CA}_{\ell}^{\mathrm{refined}}\}$
      if $k>T-\tau_{\mathrm{edit}}$ then
             //Perform edit loss update
             $\mathcal{L}_{\mathrm{edit}}\leftarrow 0$
            foreach $\mathrm{CA}_{\ell}^{\mathrm{refined}}\in\mathrm{CA}^{\mathrm{refined}}$ do
                   $\mathrm{CA}_{\mathrm{target}}\leftarrow\mathrm{ExtractTargetAttention}\Bigl{(}\mathrm{CA}_{\ell}^{\mathrm{refined}},\mathcal{I}_{\mathrm{target}}\Bigr{)}$
                  $\mathrm{CA}_{\mathrm{masked}}\leftarrow\mathrm{CA}_{\mathrm{target}}\odot\mathcal{M}$
                  $\mathcal{L}_{\mathrm{edit}}\leftarrow\mathcal{L}_{\mathrm{edit}}+\frac{\left(\sum_{(i,j)\in\mathcal{M}}\mathrm{CA}_{\mathrm{masked}}[i,j]\right)^{2}}{\sum_{(i,j)\in\mathcal{M}}\mathcal{M}[i,j]}$
            $\mathcal{L}_{\mathrm{edit}}\leftarrow\frac{\mathcal{L}_{\mathrm{edit}}}{|\mathrm{CA}^{\mathrm{refined}}|}$
            $\nabla_{\bm{z}_{\mathrm{edit}}}\mathcal{L}_{\mathrm{edit}}\leftarrow\text{Backprop}(\mathcal{L}_{\mathrm{edit}})$
            $\bm{z}_{\mathrm{edit}}\leftarrow\bm{z}_{\mathrm{edit}}-\eta\cdot\lambda_{\mathrm{edit}}\cdot\nabla_{\bm{z}_{\mathrm{edit}}}\mathcal{L}_{\mathrm{edit}}$
      //Apply classifier-free guidance during all diffusion steps
      $\bm{\epsilon}_{\mathrm{rec}}\leftarrow\mathcal{D}.\mathrm{UNet}(\bm{z}_{\mathrm{rec}},t_{k},\bm{e}_{\mathrm{rec}})$
      $\bm{\epsilon}_{\mathrm{edit}}\leftarrow\mathcal{D}.\mathrm{UNet}(\bm{z}_{\mathrm{edit}},t_{k},\bm{e}_{\mathrm{edit}},\mathrm{CA}^{\mathrm{refined}})$
      $\bm{z}_{\mathrm{rec}}\leftarrow\bm{z}_{\mathrm{rec}}-\sigma(t_{k})\cdot\bm{\epsilon}_{\mathrm{rec}}$
      $\bm{z}_{\mathrm{edit}}\leftarrow\bm{z}_{\mathrm{edit}}-\sigma(t_{k})\cdot\bm{\epsilon}_{\mathrm{edit}}$
      $\bm{z}_{\mathrm{edit}}=\mathcal{M}\cdot\bm{z}_{\mathrm{edit}}+(1-\mathcal{M})\cdot\bm{z}_{\mathrm{rec}}$
      if $k<T-\tau_{\mathrm{adv}}$ then
             //Perform pairwise adversarial optimization
            for $i\leftarrow 0$ to $max\_adv-1$ do
                   $\bm{x}_{\mathrm{rec}}\leftarrow\mathcal{D}.\mathrm{VAE}.\mathrm{Decode}(\bm{z}_{\mathrm{edit}})$
                  $\mathcal{L}_{\mathrm{adv}}\leftarrow\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}\bigl{(}\mathrm{FR}(\bm{x}_{\mathrm{rec}}),\mathrm{FR}(\tilde{\bm{x}})\bigr{)}\;+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(\bm{x}_{\mathrm{rec}},\bm{x})$
                  $\mathrm{grad}\leftarrow\text{Backprop}(\mathcal{L}_{\mathrm{adv}})$
                  $\bm{z}_{\mathrm{edit}}\leftarrow\bm{z}_{\mathrm{edit}}-\eta\cdot\mathrm{grad}$
                  $\bm{z}_{\mathrm{edit}}\leftarrow\bm{z}_{\mathrm{edit}}\odot\mathcal{M}\;+\;\bm{z}_{\mathrm{rec}}\odot\bigl{(}1-\mathcal{M}\bigr{)}$
//Decode the final latent to produce the protected image
$\bm{x}_{\mathrm{p}}\leftarrow\mathcal{D}.\mathrm{VAE}.\mathrm{Decode}(\bm{z}_{\mathrm{edit}})$
return $\bm{x}_{\mathrm{p}}$
Algorithm 1 Adversarial Makeup Generation with CA Guidance

### v2

@@
-MASQUE: A Text-Guided Diffusion-Based Framework for Localized and Customized Adversarial Makeup
+MASQUE: Localized Adversarial Makeup Generation with Text-Guided Diffusion Models
 Abstract
-As facial recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights.
+As face recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights.
 To counteract, various anti-facial recognition techniques have been proposed for privacy protection by adversarially perturbing face images, among which generative makeup-based approaches are the most popular.
@@
 Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity.
-Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity and stronger adaptability to various text makeup prompts.
+Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity preservation, stronger adaptability to various makeup prompts, and robustness to image transformations.
 1 Introduction
-Facial recognition (FR) systems [22] have been adopted in a wide range of security, biometrics, and commercial applications.
-However, their unregulated deployment poses serious privacy risks, allowing for unauthorized surveillance and malicious tracking.
-To address these concerns, anti-facial recognition (AFR) technologies [34] have emerged to protect user privacy from unauthorized FR systems.
+Facial recognition (FR) systems have been adopted in a wide range of security, biometrics, and commercial applications (Parkhi et al., 2015).
+However, their unregulated deployment poses privacy risks, allowing for unauthorized surveillance and malicious tracking.
+To address these concerns, anti-facial recognition (AFR) technologies have emerged to protect user privacy from unauthorized FR systems (Wenger et al., 2023).
 AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online.
 Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance.
-Traditional adversarial approaches, such as noise-based methods [13, 37, 41], obscure facial features with norm-bounded global perturbations, while patch-based techniques [36, 15] optimize adversarial patterns in localized image regions.
-However, these methods often introduce noticeable visual artifacts, compromising the usability of the produced images.
-To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial nets or diffusion models for AFR.
+Traditional adversarial approaches, such as noise-based methods (Joon Oh et al., 2017; Yang et al., 2021; Zhong & Deng, 2022), obscure facial features with norm-bounded global perturbations, while patch-based techniques (Xiao et al., 2021; Komkov & Petiushko, 2021) optimize adversarial patterns in localized image regions.
+Unfortunately, these methods often introduce noticeable visual artifacts, compromising the usability of the produced face images.
+To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial networks (GANs) or diffusion models for anti-facial recognition.
 Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features.
-A notable direction is makeup-based AFR [38, 12, 26, 30, 7], which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Appendix A for detailed discussions of related work).
-While makeup-based AFR methods offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts.
+A notable direction is makeup-based AFR (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025), which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Section 2 for detailed discussions of related work).
+While makeup-based AFR techniques offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts.
 In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse.
-When considering the more privacy-critical dodging scenarios, their performance in protection success rates also drops significantly.
-Contribution.
-We propose MASQUE, a novel diffusion-based image editing method for localized adversarial makeup generation with customized text guidance.
-In particular, MASQUE is designed to meet the following desiderata simultaneously:
-  • Inoffensive Identity Protection: ensure high protection success rates in dodging scenarios without requiring images of a specific target identity.
-  • Localized Modification: restrict adversarial perturbations to designated areas, preserving all other facial details.
-  • User Control: demonstrate strong prompt-following ability and adaptability to diverse text makeup prompts.
+When considering the more privacy-critical dodging scenarios, their performance in protection success rates often drops significantly.
+Table 1: Highlights of the key desirable features of MASQUE compared with prior AFR methods.
+Dodging External ID Guidance Localization Prompt-Following Ability
+TIP-IM (Yang et al., 2021) - -
+AMT-GAN (Hu et al., 2022) image low
+C2P (Shamshad et al., 2023) text medium
+DiffAM (Sun et al., 2024) image medium
+MASQUE (ours) text high
+Contributions.
+Recognizing a few limitations of prior state-of-the-art AFR methods (Section 4.1), we develop MASQUE, a novel diffusion-based image editing framework for localized adversarial makeup generation with customized text guidance (Section 4.2).
+In particular, MASQUE stands out by achieving the following desiderata simultaneously (see Table 1 for a summary of its key distinctive features):
+  • Inoffensive Identity Protection.
+Built upon pairwise adversarial guidance, MASQUE ensures high dodging success rates without requiring face images of any external identity other than the victim’s.
+Such a property reduces the risks of targeted misuse and ethical concerns (Section 3.3).
+  • Localized Modification.
+MASQUE employs a facial mask generation and regularization module to constrain adversarial perturbations to designated areas, while preserving fine-grained details of the original face image, thereby vastly enhancing the visual quality of protected images.
+  • Strong User Control.
+By utilizing cross-attention fusion with masking, MASQUE achieves strong prompt-following capability and is adaptable to diverse makeup prompts, providing greater user control and convenience than existing state-of-the-art AFR methods.
+Through comprehensive experiments on two popular image benchmarks, a range of FR models and their variants, along with multiple makeup text prompts, MASQUE not only shows significantly higher dodging success rates over baseline methods, but is also capable of preserving visual quality and fine details of original face images (Section 5).
+In addition, we also demonstrate the robustness of MASQUE’s performance to diverse makeup prompts and various image transformations (Section 6).
+All of the above suggest the potential of MASQUE as a superior solution to real-world AFR applications.
+2 Related Work
+Anti-Facial Recognition.
+Earlier works adopted obfuscation techniques to obscure the facial identity features (Newton et al., 2005) or craft $\ell_{p}$ perturbations to fool FR models (Yang et al., 2021).
+While effective, these methods often compromise image quality, limiting their real-world applicability.
+Poisoning-based approaches (Cherepanova et al., 2021; Shan et al., 2020) introduce a new protection scheme by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models.
+These techniques excel at disrupting model training without visibly altering the image, but rely on strong assumptions about when and how the unauthorized FR model is constructed.
+Recently, adversarial makeup (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025) has emerged as a promising solution to realize the goal of facial privacy protection.
+These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while preserving the aesthetic appeal.
+For example, Hu et al.
+(2022) proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the GAN framework (Goodfellow et al., 2014).
+Advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect (C2P) (Shamshad et al., 2023) and DiffAM (Sun et al., 2024), which adopt a StyleGAN model or a diffusion-based framework, to enable seamless adversarial face modifications with much improved visual quality.
+In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for facial privacy protection under dodging scenarios.
+More recently, a line of research (Liu et al., 2023; 2024; Wang et al., 2025; Han et al., 2025; Salar et al., 2025) proposed leveraging diffusion models to generate imperceptible AFR-protections without focusing on makeups.
+For instance, DiffProtect (Liu et al., 2023) uses the diffusion autoencoder model and adversarially guides the semantic code of the original image towards a target face, whereas Salar et al.
+(2025) improved the prior approach by learning unconditional embeddings as null-text guidance and adversarially modifying the latent code in the latent diffusion model.
+Our work complements these methods, where we focus on improving the state-of-the-art makeup-based AFR approaches without relying on images from any external identity, while preserving the visual quality of original face images with a strong emphasis on localizing and controlling the added perturbations.
+Diffusion-Based Image Editing.
+Localized editing methods extend diffusion models for targeted modifications.
+For example, mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency.
+Attention-based approaches (Hertz et al., 2022; Tumanyan et al., 2023) guide edits via attention injection, achieving better global structure preservation but suffering from editing leakage.
+In addition, text-guided diffusion models, capable of synthesizing high-quality images from natural language descriptions, significantly promote the popularization of generative AI.
+Built on denoising diffusion probabilistic models (Ho et al., 2020), they iteratively refine random noise into coherent images guided by text prompts.
+Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation.
+Mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions.
+While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios.
+On the other hand, mask-free attention-based methods (Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023) use attention injection mechanisms to guide edits without requiring explicit masks.
+These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region.
+In this work, we leverage mask-based cross-attention guidance (Mao et al., 2023) to achieve localized adversarial perturbations in the form of makeup.
+This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts, while addressing the limitations of structural inconsistencies seen in previous methods.
 3 Problem Formulation
-In this section, we specify the threat model for unauthorized facial recognition (Section 2.1), introduce the background on anti-facial recognition (Section 2.2), and explain why dodging is more suitable than impersonation (Section 2.3).
-2.1 Threat Model
-We consider the adversarial setup, where an attacker adopts unauthorized FR models to identify benign users from their publicly shared facial images.
-This enables the extraction of sensitive information, posing significant privacy risks.
-Since many well-trained FR models are readily available—either as open-source implementations or through commercial APIs—attackers can easily obtain automated FR tools to achieve this malicious objective.
-Formally, let $\mathcal{X}$ be the input facial image space and $\mathcal{Y}$ be the set of possible identities.
-Given a collection of online-scraped face images $\mathcal{D}$, the adversary aims to correctly identify victim users from their facial images according to the following objective:
+3.1 Facial Recognition (FR)
 […]
-where $x$ stands for a face image, $y$ is the corresponding ground-truth identity, and $\mathrm{FR}:\mathcal{X}\rightarrow\mathcal{Y}$ denotes a FR model.
-A typical FR model consists of three key components: a feature extractor, a gallery database, and a query matching step (see [34, 7] for detailed descriptions).
-Aligned with prior work [26, 30], we assume that FR models employed by the attacker can achieve high accuracy on clean images.
-Since FR systems can vary in design, the attacker may utilize different feature extractors and query-matching schemes.
+Typically, a FR model consists of three components: a feature extractor, a gallery database, and a query matching scheme (see Wenger et al.
+(2023) for detailed descriptions).
+Since FR models can vary in design, the attacker may utilize different feature extractors and query-matching schemes.
 To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs.
-Additionally, we consider face verification as a specialized FR variant, where the gallery database contains only a single user.
-2.2 Anti-Facial Recognition
-The primary goal of AFR is to evade the identification of unauthorized facial recognition models used by adversaries described in Section 2.1, improving the protection of benign users’ privacy.
-In this work, we focus on the most popular adversarial-based AFR techniques, which craft imperceptible or naturalistic perturbations (e.g., adversarial makeup) to user’s images to fool FR systems.
-Specifically, the objective of adversarial-based AFR can be cast into a constrained optimization problem as follows:
-$\max\>\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}(x)\big{)}\neq y\big{\}}$ (2)
-[…]
-where $\mathrm{AFR}:\mathcal{X}\rightarrow\mathcal{X}$ denotes the AFR perturbation function, $\Delta$ is a similarity metric that measures the visual distortion of the AFR-perturbed image $\mathrm{AFR}(x)$ with reference to the original image $x$, and $\gamma>0$ is a threshold parameter reflecting the distortion upper bound the user can tolerate.
-Dodging Success.
-We consider benign users who seek privacy protection through adversarial-based AFR before sharing their facial images online.
-A desirable AFR technique should achieve a high dodging success rate (DSR) against unauthorized facial recognition, as defined by the optimization objective in Equation LABEL:eq:adversarial_AFR.
-For clarity, we refer to these users as defenders, as they apply AFR techniques to modify their facial images with adversarial perturbations.
+In addition, we consider face verification as a specialized FR variant, where the gallery database contains only a single user’s images.
+3.2 Anti-Facial Recognition (AFR)
+The goal of AFR is to evade face recognition, thereby improving the protection of benign users’ privacy.
+We focus on the most popular adversarial-based AFR techniques, which craft imperceptible or naturalistic perturbations (e.g., adversarial makeup) to users’ face images to fool FR models.
+Specifically, the objective can be cast into a constrained optimization problem detailed as follows:
+$\max\>\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}(\bm{x})\big{)}\neq y\big{\}},\quad\text{s.t.}\>\>\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\Delta\big{(}\mathrm{AFR}(\bm{x}),\bm{x}\big{)}\leq\gamma,$ (1)
+where $\mathrm{AFR}:\mathcal{X}\rightarrow\mathcal{X}$ denotes the AFR perturbation function, $\Delta$ is a similarity metric that measures the visual distortion of the AFR-perturbed image $\mathrm{AFR}(\bm{x})$ with reference to the original image $\bm{x}$, and $\gamma>0$ is a threshold parameter reflecting the distortion upper bound that can be tolerated.
+As characterized by the optimization objective in Equation 1, a desirable AFR technique is expected to achieve a high dodging success rate (DSR) against facial recognition, which is the primary evaluation metric when comparing the effectiveness of different AFR techniques.
+For simplicity, we also refer to these users as defenders, who apply AFR techniques to modify their facial images with adversarial perturbations.
 Since the defender typically lacks precise knowledge of the FR model employed by the attacker, we consider the black-box scenario when evaluating the DSR of each AFR method.
-Visual Quality.
-From the user’s perspective, images perturbed by AFR should appear natural and closely resemble the original, as captured by the optimization constraint in Equation LABEL:eq:adversarial_AFR.
-If $\mathrm{AFR}(x)$ has an unsatisfactory image quality, users may be reluctant to share the protected face image online, even if a high DSR is achieved.
-Given this expectation of high visual quality, earlier privacy-preserving techniques such as face obfuscation or anonymization [20, 29] are not suitable.
-In comparison, makeup-based AFR methods [26, 30] stand out, as they introduce natural adversarial perturbations with minimal visual distortion.
-In this work, we further improve the performance of generative makeup-based AFR, enhancing both DSR and visual quality.
-Figure 1: Each row presents the reference image with a text makeup prompt, the original image, followed by the generations of each AFR method.
-Methods marked with $\dagger$ require only a text prompt for makeup application and do not rely on a reference image.
-The yellow text indicates the cosine similarity score.
-Below the images, we compare ISR and DSR averaged across four FR models and three makeup styles.
+From the defender’s perspective, images perturbed by AFR should appear natural and closely resemble the original, as captured by the optimization constraint in Equation 1.
+If $\mathrm{AFR}(\bm{x})$ has an unsatisfactory image quality, users may be reluctant to share the protected face image online, even if a high DSR is achieved.
+Given this expectation of high visual quality, earlier privacy-preserving techniques such as face obfuscation or anonymization (Newton et al., 2005; Sun et al., 2018) are not suitable.
+In comparison, makeup-based AFR methods (Shamshad et al., 2023; Sun et al., 2024) stand out, as they introduce natural adversarial perturbations with minimal visual distortion.
+In this work, we aim to improve the performance of generative makeup-based AFR, enhancing both DSR and visual quality.
 3.3 Why Dodging instead of Impersonation?
-In contrast to the untargeted dodging objective, several state-of-the-art studies [37, 12, 26, 30] have focused on improving AFR performance in targeted impersonation scenarios.
+In contrast to the untargeted dodging objective, several state-of-the-art studies (Yang et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024) have focused on improving AFR performance in targeted impersonation scenarios.
 Impersonation-oriented methods aim to adversarially modify a user’s facial images to mislead a FR model into recognizing them as a specific target identity.
-Let $\mathrm{AFR}_{\mathrm{I}}:\mathcal{X}\rightarrow\mathcal{X}$ be the perturbation function for an impersonation-oriented method.
-In existing literature, the impersonation success rate (ISR) is typically measured by the following metric:
-$\text{ISR}:=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}_{\mathrm{I}}(x)\big{)}=y_{t}\},$ (3)
-where $y_{t}$ denotes the target identity satisfying $y_{t}\neq y$. Below, we further detail why impersonation is not suitable for AFR.
-Misaligned Goal.
-From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics, ISR and DSR.
-As shown by [42], high ISR does not guarantee high DSR.
+Let $\mathrm{AFR}_{\mathrm{I}}:\mathcal{X}\rightarrow\mathcal{X}$ be the perturbation function with respect to an impersonation-oriented method.
+In prior work, the impersonation success rate (ISR) is usually computed using the following metric:
+$\text{ISR}:=\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}_{\mathrm{I}}(\bm{x})\big{)}=y_{t}\},$ (2)
+where $y_{t}$ is the target identity with $y_{t}\neq y$. Below, we explain why impersonation is not appropriate.
+From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics (ISR and DSR).
+As shown by Zhou et al.
+(2024), high ISR does not guarantee high DSR.
 If the target identity $y_{t}$ is not included in the attacker’s model (e.g., using facial verification with only a single image of the victim identity $y$), optimizing for higher ISR offers little benefit due to misaligned objectives.
-As we will illustrate in Section 3.1, prior impersonation-oriented AFR methods can only achieve weak privacy protection when evaluated under dodging scenarios.
-Therefore, we argue that effective privacy protection should prioritize untargeted dodging success.
-Targeted Misuse.
-Impersonation increases the risk of targeted misuse, raising serious ethical and fairness concerns.
-Unlike dodging, impersonation-based approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse.
+As will be shown in Section 1, prior impersonation-oriented AFR methods can only achieve weak protection performance when evaluated in dodging scenarios.
+In addition, impersonation techniques can lead to a higher likelihood of targeted misuse.
+Unlike dodging, impersonation-oriented approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse.
 This ability can be exploited for malicious purposes, such as falsely associating one’s actions with another individual.
 For instance, if a user $y$ intentionally applies an AFR tool to impersonate a target $y_{t}$ and engages in illegal activity, FR systems may misidentify $y_{t}$ as the suspect.
-This deliberate identity manipulation poses significant ethical and legal risks [34].
-By contrast, dodging-oriented AFR serves as a better solution for facial privacy protection.
+This deliberate identity manipulation poses significant ethical and legal risks (Wenger et al., 2023).
+Therefore, effective privacy protection should aim for untargeted dodging success.
 Instead of redirecting recognition to a specifically chosen identity, dodging causes misclassification in an uncontrolled manner, typically as an unknown or incorrect identity.
@@
 While dodging still prevents correct identification, it does so without assigning a specific false identity, reducing the risk of targeted deception or malicious intent.
-Dodging External ID Guidance Local Prompt-Following
-TIP-IM [37] n/a n/a
-AMT-GAN [12] image low
-C2P [26] text medium
-DiffAM [30] image medium
-MASQUE (ours) text high
-Table 1: Comparisons of SOTA AFR methods for key features in terms of both privacy protection and visual quality.
-3 Challenges of Existing AFR Methods
-In this section, we present preliminary results to highlight key limitations of previous work, further motivating the design of our approach, which will be detailed in Section 4.
-Preliminary Results.
-Figure 1 shows qualitative comparisons of three makeup styles, evaluating several prior AFR methods, including a noise-based method, TIP-IM [37], and three state-of-the-art makeup-based approaches [12, 26, 30].
-For each AFR method, we report the metrics across different black-box facial verification models averaged across $100$ randomly selected images from CelebA-HQ [14].
-To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset [16] or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 for more details).
-Table 1 summarizes the key feature comparisons of state-of-the-art AFR methods, including ours.
-3.1 Weak Protection under Dodging
-As discussed in Section 2.3, an ideal AFR should account for dodging scenarios, to ensure privacy without dependence on a target identity.
+4 Proposed Method: MASQUE
+4.1 Limitations of Existing AFR Methods
+Figure 1: Each row presents the reference image with a text makeup prompt, the original image, followed by AFR-protected images.
+Methods marked with $\dagger$ require only a text prompt for makeup application and do not rely on a reference image.
+The yellow text indicates the cosine similarity score.
+Below the images, we report the averaged impersonation and dodging success rates of each method.
+We conduct preliminary experiments to understand the performance of existing AFR methods in terms of privacy protection and visual quality.
+Figure 1 shows qualitative comparisons of three makeup styles (“red lipstick”, “blue eyebrow”, and “pink eyeshadow”), evaluating several adversarial-based AFR methods, including one of the best current noise-based methods, TIP-IM (Yang et al., 2021), and three state-of-the-art makeup-based approaches (Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024).
+For each method, we report both the metrics of impersonation and doding success rates across four black-box facial verification models averaged across 100100100 randomly selected images from CelebA-HQ (Karras et al., 2017) and the three makeup styles.
+To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset (Li et al., 2018) or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 and Appendix B for more experimental details of this preliminary study).
+Weak Protection under Dodging.
+As discussed in Section 3.3, an ideal AFR should account for dodging scenarios to ensure privacy without dependence on a target identity.
 However, despite its ethical advantages, dodging remains an underexplored and underdeveloped area in makeup-based AFR.
@@
 This highlights a fundamental gap: current techniques are not optimized for dodging, leaving privacy protection less reliable when a user seeks to avoid recognition rather than impersonate another identity.
-In addition, while some methods like [26, 7] introduce a form of dodging, they still rely on a target identity.
+In addition, while some AFR methods like Shamshad et al.
+(2023); Fan et al.
+(2025) introduce a variant for dodging, they still rely on a target identity.
 This reliance contradicts the very goal of dodging-based AFR, which should function without any identity substitution.
 Thus, a key challenge in AFR research is developing a robust dodging-based approach that provides strong protection while avoiding the ethical risks associated with impersonation.
-3.2 Non-Localized Face Modification
-Noise-based approaches like TIP-IM generate global perturbations, which fail to localize the editings.
-Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches, including AdvMakeup, AMT-GAN, and C2P, introduce global artifacts that alter non-facial regions.
+Non-Localized Face Modification.
+Noise-based approaches (TIP-IM) generate global perturbations, which fail to localize the edits.
+Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches (AdvMakeup, AMT-GAN, and C2P) introduce global artifacts that alter non-facial regions.
 This lack of precise localization is particularly problematic when users seek to protect only their facial identity while preserving the background.
 As shown in Figure 1, this issue is especially pronounced in complex backgrounds, where perturbations should remain confined to facial regions but instead extend beyond.
-In contrast, the diffusion-based method DiffAM enables better localization, effectively applying makeup perturbations from a reference image.
+In contrast, DiffAM enables better localization, effectively applying makeup perturbations from a reference image.
 However, its reliance on reference images limits flexibility, particularly when suitable references are unavailable.
 This highlights the need for methods that achieve both precise localization and greater flexibility.
-3.3 Limited Prompt-Following Ability
-Makeup-based AFR methods aim to modify facial appearance to prevent accurate identification by FR systems.
-However, existing methods offer limited user control, making customization and usability a key challenge.
-Broadly, makeup-based AFR methods fall into two categories:
+Limited Prompt-Following Ability.
+Makeup-based AFR aim methods aim to modify facial appearance to prevent accurate identification by FR models.
+However, existing methods offer limited user control, making customization and usability a challenge.
+Below, we elaborate on this challenge for makeup-based AFR methods based on reference images and text prompts, respectively:
   • Image-reference-based AFR transfers makeup styles from a reference image.
@@
 It only applies makeup to three fixed regions—skin, mouth, and eyes—preventing transfer to other areas like colored eyebrows.
-Additionally, these methods require separate fine-tuning for each reference, limiting efficiency.
+In addition, these methods require separate fine-tuning for each reference image, incurring high computational overhead.
   • Text-prompt-based AFR allows users to describe makeup styles in natural language, eliminating the need for reference images or separate fine-tuning while enhancing privacy.
 However, this method lacks fine-grained control, as C2P relies solely on CLIP directional loss, often leading to unintended modifications beyond the intended areas.
-4 Our Methodology: MASQUE
+4.2 Detailed Design
 Figure 2: The pipeline of MASQUE involves: (1) fusing the editing and reconstruction prompts to produce an updated cross-attention map for diffusion, (2) creating a mask $\mathcal{M}$ to define a target region and optimize an edit loss to maximize makeup-related attention in $\mathcal{M}$, and (3) using pairwise adversarial guidance with same-identity image to enhance identity confusion without external targets.
-To address the key challenges discussed in Section 3, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced.
-Figure 2 illustrates the pipeline of MASQUE with its pseudocode in Appendix D.
-The makeup is guided by a user-defined text prompt $p^{*}$, refined via cross-attention fusion with a mask.
-Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR systems without compromising visual fidelity.
-Our method leverages the state-of-the-art Stable Diffusion [24] adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent $z_{T}$ conditioned on a text embedding $C$.
-To be more specific, Stable Diffusion is trained to predict the added noise $\epsilon$ via:
-$\min_{\theta}\ \mathbb{E}_{{z}_{0},\epsilon\sim\mathcal{N}(0,I),t\sim\mathrm{Unif}(1,T)}\bigg{[}\big{\|}\epsilon-\epsilon_{\theta}(z_{t},t,C)\big{\|}_{2}^{2}\bigg{]}.$
+To address the limitations of existing AFR methods discussed in Section 4.1, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced.
+Figure 2 illustrates the pipeline of MASQUE with its pseudocode detailed in Algorithm 1 in Appendix A.
+The makeup generation process is guided by a user-defined text prompt $p^{*}$, refined via cross-attention fusion with a mask.
+Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR models without compromising visual fidelity.
+In particular, our method leverages the state-of-the-art Stable Diffusion (Rombach et al., 2022) adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent ${\bm{z}}_{T}$ conditioned on a text embedding $\bm{c}$.
+To be more specific, Stable Diffusion is trained to predict the added noise $\bm{\epsilon}$ via:
+$\min_{\theta}\ \mathbb{E}_{{\bm{z}}_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathrm{I}),t\sim\mathrm{Unif}(1,T)}\bigg{[}\big{\|}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})\big{\|}_{2}^{2}\bigg{]}.$ (3)
 Applying Makeup with Cross Attention.
-Before applying edits, we first obtain a faithful latent representation of the original image $x$ using null-text inversion [19], which mitigates reconstruction errors common in direct DDIM inversion [28].
+Before applying edits, we first obtain a faithful latent representation of the original image $\bm{x}$ using null-text inversion (Mokady et al., 2023), which mitigates reconstruction errors common in direct DDIM inversion (Song et al., 2020).
 Conditioning on an empty prompt to align the forward and reverse diffusion trajectories, ensuring near-perfect reconstruction of the image’s structure and identity.
-With this accurate latent representation, we introduce makeup attributes via the text prompt $p^{*}$ by manipulating the cross-attention (CA) layers of the diffusion model [24], which control how spatial features correspond to semantic tokens.
+With this accurate latent representation, we introduce makeup attributes via the text prompt $p^{*}$ by manipulating the cross-attention (CA) layers of the diffusion model (Rombach et al., 2022), which control how spatial features correspond to semantic tokens.
 At each diffusion step $\tau$, we extract attention maps $A_{\tau}$ (reconstruction) and $A_{\tau}^{*}$ (editing) and blend them: preserving CA values from $A_{\tau}$ for shared tokens to maintain structure, while incorporating values from $A_{\tau}^{*}$ for makeup-specific tokens in $p^{*}$:
-$\big{(}\mathrm{Update}(A_{\tau},A_{\tau}^{*})\big{)}_{i,j}:=\begin{cases}(A_{\tau})_{i,j},&\text{if $j$ is in both $p$ and $p^{*}$},\\[6.0pt] (A_{\tau}^{*})_{i,j},&\text{if $j$ is unique to $p^{*}$},\end{cases}$
-where $p$ denotes the original text prompt. The result is $\hat{A}_{\tau}$, a set of mixed CA maps that preserve the original facial layout while steadily introducing adversarial makeup features [9].
+[…]
+where $p$ denotes the original text prompt.
+The result is $\hat{A}_{\tau}$, a set of mixed CA maps that preserve the original facial layout while steadily introducing adversarial makeup features (Hertz et al., 2022).
 Enhancing Semantic Edits and Locality.
 To ensure precise localization, we generate a mask $\mathcal{M}$ that defines the region for modification.
-To achieve this, we embed the prompt $p^{*}$ using a Sentence Transformer [23] model and compare it to embeddings of predefined facial regions.
+To achieve this, we embed the prompt $p^{*}$ using a Sentence Transformer (Reimers & Gurevych, 2019) model and compare it to embeddings of predefined facial regions.
 The closest match determines the relevant area for the edit.
@@
 Once the target region $\mathcal{M}$ is determined, we enhance the influence of makeup-related tokens by maximizing their attention within $\mathcal{M}$.
-This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality [18].
+This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality (Mao et al., 2023).
 Specifically, we optimize:
-$\mathcal{L}_{\mathrm{edit}}=\left(1-\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\frac{(A_{\tau}^{*})_{i,\mathrm{new}}}{(A_{\tau}^{*})_{i,\mathrm{new}}+\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}}\right)^{2},$
+$\mathcal{L}_{\mathrm{edit}}=\left(1-\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\frac{(A_{\tau}^{*})_{i,\mathrm{new}}}{(A_{\tau}^{*})_{i,\mathrm{new}}+\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}}\right)^{2},$ (5)
 where $(A_{\tau}^{*})_{i,\mathrm{new}}$ is the attention weights assigned to the new makeup tokens at spatial index $i$, while $\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}$ is the total attention weight of tokens that appear in both the original and makeup prompts.
-By emphasizing makeup-specific attention in the masked region, this loss term ensures that modifications in $x_{\mathrm{p}}$ occur meaningfully in the intended region while preventing edits from spreading elsewhere.
-To further enforce spatial precision, we process the latent through two distinct branches: an edit branch for makeup application and a reconstruction branch to preserve original features.
-During the backward steps, we explicitly constrain perturbations within the designated area by imposing the following conditions:
-$z_{\text{edit}}=\mathcal{M}\cdot z_{\text{edit}}+(1-\mathcal{M})\cdot z_{\text{rec}},$
-where $z_{\text{edit}}$ represents the latent from the edit branch, and $z_{\text{rec}}$ denotes the latent from the reconstruction branch.
+By emphasizing makeup-specific attention in the masked region, this loss term ensures that modifications in $\bm{x}_{\mathrm{p}}$ occur meaningfully in the intended region while preventing edits from spreading elsewhere.
+To enforce spatial precision, we process latents through two distinct branches: an edit branch for makeup application and a reconstruction branch to preserve original features.
+During the backward steps, we explicitly constrain perturbations within the designated area by imposing:
+$\bm{z}_{\text{edit}}=\mathcal{M}\cdot\bm{z}_{\text{edit}}+(1-\mathcal{M})\cdot\bm{z}_{\text{rec}},$ (6)
+where $\bm{z}_{\text{edit}}$ denotes the latent from the edit branch, and $\bm{z}_{\text{rec}}$ is the latent from the reconstruction branch.
 While makeup edits ensure semantic plausibility, the core adversarial objective is to disrupt FR models.
-Here lies our novel contribution: a pairwise adversarial guidance that uses a guide image of the same identity to achieve robust identity confusion without relying on an external target.
+Therefore, we propose a novel pairwise adversarial guidance loss that uses a guide image of the same identity to achieve robust identity confusion without relying on an external target.
+Pariwise Adversarial Guidance.
+Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios.
+By contrast, our approach leverages a pair $(\bm{x},\tilde{\bm{x}})$ of face images from the same individual, where $\tilde{\bm{x}}$ serves as the guide image.
+This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss.
+Note that in the standard diffusion denoising process, the objective is to generate images similar to the original.
+If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness.
+In our evaluations, we test the impact of the number of guide images $G$ on MASQUE’s performance (Appendix C.1) and whether using self-augmentation without a separate guide image ($G=0$) in MASQUE can be an alternative (Appendix C.2).
+Adversarial and Image Quality Trade-off.
+We introduce adversarial perturbations during the later stage of the diffusion process, ensuring the coarse structure remains intact while subtly altering identity-specific features.
+To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure:
+$\mathcal{L}_{\mathrm{adv}}=\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}(\bm{x}_{\mathrm{p}},\tilde{\bm{x}})+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(\bm{x}_{\mathrm{p}},\bm{x}),$ (7)
+where $\bm{x}_{\mathrm{p}},\tilde{\bm{x}}$ refer to the protected and guide images, and $\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$ are trade-off parameters.
+Here, $\mathrm{CoSi}(\bm{x}_{\mathrm{p}},\tilde{\bm{x}})$ ensures that the perturbations sufficiently diverge from recognizable identity features, while the LPIPS loss (Zhang et al., 2018) maintains perceptual and structural fidelity.
+5 Experiments
+In this section, we conduct experiments to benchmark the performance of MASQUE in dodging success and visual aspects, with comparisons to state-of-the-art makeup-based AFR methods.
+Due to space limits, we present additional experiments, ablation studies, and visualizations in Appendices B-E.
+Table 2: DSR (%) of various AFR methods under facial identification and verification settings. Here, $G$ denotes the number of guide images used in MASQUE. Each best result is highlighted in bold.
 Method CelebA-HQ VGG-Face2-HQ Avg.
 IR152 IRSE50 FaceNet MobileFace IR152 IRSE50 FaceNet MobileFace
-Identification Clean $10.00$ $13.00$ $5.00$ $40.00$ $13.00$ $18.00$ $18.00$ $25.00$ $17.75$
-TIP-IM $62.00$ $86.00$ $65.00$ $74.00$ $57.00$ $73.00$ $52.00$ $62.00$ $66.38$
-AMT-GAN $61.22$ $48.33$ $50.67$ $57.33$ $16.00$ $24.00$ $22.33$ $32.33$ $39.03$
-DiffAM $54.00$ $57.67$ $59.00$ $74.33$ $49.67$ $51.67$ $54.00$ $70.67$ $58.88$
-C2P (I) $30.67$ $38.33$ $22.00$ $56.67$ $18.00$ $20.00$ $21.67$ $32.00$ $29.92$
-C2P (D) $74.67$ $76.33$ $56.33$ $77.33$ $18.33$ $19.67$ $20.33$ $30.33$ $46.67$
-MASQUE ($G=0$) $92.33$ $95.67$ $61.33$ $87.00$ $69.67$ $76.33$ $45.33$ $81.67$ $76.17$
-MASQUE ($G=1$) $\underline{98.00}$ $\underline{98.33}$ $\underline{77.67}$ $\underline{94.00}$ $\underline{84.33}$ $\underline{89.00}$ $\underline{61.00}$ $\underline{91.00}$ $\underline{85.50}$
-Verification Clean $5.00$ $5.00$ $4.00$ $10.00$ $13.00$ $17.00$ $15.00$ $30.00$ $12.38$
-TIP-IM $44.00$ $60.00$ $51.00$ $40.00$ $46.00$ $51.00$ $59.00$ $46.00$ $49.63$
-AMT-GAN $40.00$ $32.00$ $51.67$ $25.67$ $14.00$ $26.33$ $26.67$ $40.67$ $32.13$
-DiffAM $31.00$ $26.00$ $54.00$ $43.33$ $46.33$ $50.67$ $56.00$ $82.67$ $48.75$
-C2P (I) $12.33$ $11.33$ $14.33$ $21.33$ $16.33$ $18.67$ $19.33$ $37.33$ $18.88$
-C2P (D) $52.33$ $46.67$ $40.33$ $41.67$ $17.00$ $19.67$ $18.33$ $38.00$ $34.25$
-MASQUE ($G=0$) $89.33$ $90.67$ $57.67$ $76.33$ $62.33$ $71.33$ $42.33$ $86.00$ $72.00$
-MASQUE ($G=1$) $\underline{96.00}$ $\underline{96.67}$ $\underline{75.67}$ $\underline{79.33}$ $\underline{82.33}$ $\underline{91.67}$ $\underline{63.33}$ $\underline{94.33}$ $\underline{84.92}$
-Table 2: Dodging success rate (%) under black-box facial identification and verification settings.
-For each column, the other three FR systems are used as surrogates to generate the protected faces, and the best result is underlined.
-Here, $G$ denotes the number of guide images.
-Pariwise Adversarial Guidance.
-Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios.
-By contrast, our approach leverages a pair $(x,\tilde{x})$ of face images from the same individual, where $\tilde{x}$ serves as the guide image.
-This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss.
-Note that in the standard diffusion denoising process, the objective is to generate images similar to the original.
-If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness (see Tables 2-3 in Appendix B for supporting evidence and comparison results with self-augmented guide images).
-Adversarial and Image Quality Trade-off.
-We introduce adversarial perturbations during the later stage of the diffusion process, ensuring coarse structure remains intact while subtly altering identity-specific features.
-To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure that captures adversarial effectiveness:
-$\mathcal{L}_{\mathrm{adv}}=\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}(x_{\mathrm{p}},\tilde{x})+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(x_{\mathrm{p}},x),$
-where $x_{\mathrm{p}}$ denotes the protected image.
-$\mathrm{CoSi}$ ensures that the adversarial perturbations sufficiently diverge from recognizable identity features, while $\mathrm{LPIPS}$ [39] maintains perceptual and structural fidelity, respectively.
-The parameters $\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$ allow fine-tuning of this trade-off.
-We conduct ablation studies to understand the sensitivity of these hyperparameters and how to set them in MASQUE for better trade-offs (see Section 5.3 and Appendix B).
-5 Experiments
-In this section, we conduct comprehensive experiments to study the effectiveness of MASQUE with comparisons to prior AFR methods in both privacy protection and visual aspects.
-Appendices B and C present additional experimental results and visualizations for more general makeup prompts.
-5.1 Experimental Settings
+Identification Clean 10.0010.0010.00 13.0013.0013.00 5.005.005.00 40.0040.0040.00 13.0013.0013.00 18.0018.0018.00 18.0018.0018.00 25.0025.0025.00 17.7517.7517.75
+TIP-IM 62.0062.0062.00 86.0086.0086.00 65.0065.0065.00 74.0074.0074.00 57.0057.0057.00 73.0073.0073.00 52.0052.0052.00 62.0062.0062.00 66.3866.3866.38
+AMT-GAN 61.2261.2261.22 48.3348.3348.33 50.6750.6750.67 57.3357.3357.33 16.0016.0016.00 24.0024.0024.00 22.3322.3322.33 32.3332.3332.33 39.0339.0339.03
+DiffAM 54.0054.0054.00 57.6757.6757.67 59.0059.0059.00 74.3374.3374.33 49.6749.6749.67 51.6751.6751.67 54.0054.0054.00 70.6770.6770.67 58.8858.8858.88
+C2P (I) 30.6730.6730.67 38.3338.3338.33 22.0022.0022.00 56.6756.6756.67 18.0018.0018.00 20.0020.0020.00 21.6721.6721.67 32.0032.0032.00 29.9229.9229.92
+C2P (D) 74.6774.6774.67 76.3376.3376.33 56.3356.3356.33 77.3377.3377.33 18.3318.3318.33 19.6719.6719.67 20.3320.3320.33 30.3330.3330.33 46.6746.6746.67
+MASQUE ($G=0$) 92.3392.3392.33 95.6795.6795.67 61.3361.3361.33 87.0087.0087.00 69.6769.6769.67 76.3376.3376.33 45.3345.3345.33 81.6781.6781.67 76.1776.1776.17
+MASQUE ($G=1$) $\bm{98.00}$ $\bm{98.33}$ $\bm{77.67}$ $\bm{94.00}$ $\bm{84.33}$ $\bm{89.00}$ $\bm{61.00}$ $\bm{91.00}$ $\bm{85.50}$
+Verification Clean 5.005.005.00 5.005.005.00 4.004.004.00 10.0010.0010.00 13.0013.0013.00 17.0017.0017.00 15.0015.0015.00 30.0030.0030.00 12.3812.3812.38
+TIP-IM 44.0044.0044.00 60.0060.0060.00 51.0051.0051.00 40.0040.0040.00 46.0046.0046.00 51.0051.0051.00 59.0059.0059.00 46.0046.0046.00 49.6349.6349.63
+AMT-GAN 40.0040.0040.00 32.0032.0032.00 51.6751.6751.67 25.6725.6725.67 14.0014.0014.00 26.3326.3326.33 26.6726.6726.67 40.6740.6740.67 32.1332.1332.13
+DiffAM 31.0031.0031.00 26.0026.0026.00 54.0054.0054.00 43.3343.3343.33 46.3346.3346.33 50.6750.6750.67 56.0056.0056.00 82.6782.6782.67 48.7548.7548.75
+C2P (I) 12.3312.3312.33 11.3311.3311.33 14.3314.3314.33 21.3321.3321.33 16.3316.3316.33 18.6718.6718.67 19.3319.3319.33 37.3337.3337.33 18.8818.8818.88
+C2P (D) 52.3352.3352.33 46.6746.6746.67 40.3340.3340.33 41.6741.6741.67 17.0017.0017.00 19.6719.6719.67 18.3318.3318.33 38.0038.0038.00 34.2534.2534.25
+MASQUE ($G=0$) 89.3389.3389.33 90.6790.6790.67 57.6757.6757.67 76.3376.3376.33 62.3362.3362.33 71.3371.3371.33 42.3342.3342.33 86.0086.0086.00 72.0072.0072.00
+MASQUE ($G=1$) $\bm{96.00}$ $\bm{96.67}$ $\bm{75.67}$ $\bm{79.33}$ $\bm{82.33}$ $\bm{91.67}$ $\bm{63.33}$ $\bm{94.33}$ $\bm{84.92}$
+5.1 Experimental Setup
 Dataset.
-To reflect the high-quality nature of online facial images, we conduct experiments using $300$ images from each of the CelebA-HQ [14] and VGG-Face2-HQ [3] datasets, all at a resolution of $1024\times 1024$.
-We randomly sample $100$ identities from per dataset, each with three images: a probe image to be protected, a reference image for guidance, and a the gallery image, which is assumed to be stored in the facial recognition system for comparison.
+To reflect the high-quality nature of online face images, we use 300300300 high-resolution images ($\times 1024$ 1024) from each of the CelebA-HQ (Karras et al., 2017) and VGG-Face2-HQ (Chen et al., 2024) datasets.
+For each dataset, we randomly sample 100 identities, each with three images: a probe $\bm{x}$ (to be protected), a reference $\tilde{\bm{x}}$ (as the guide image), and a gallery image $\bm{x}_{\mathrm{g}}$ (stored in the attacker’s facial recognition model).
+Note that the VGG-Face2-HQ dataset contains face images under challenging conditions, such as pose variation and natural occlusions like sunglasses and hair.
 Configuration.
-We compare our methods against the following baselines: TIP-IM [37], the SOTA noised-based AFR, and other recent generative-based AFR methods, including AMT-GAN [12], Clip2Protect [26], and DiffAM [30].
-For Clip2Protect, we denote it as C2P for simplicity, where both the impersonating version (denoted as C2P (I)) and the dodging version (i.e., C2P (D)) are assessed.
-We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 [6], IRSE50 [11], FaceNet [25], and MobileFace [2].
-We also evaluate AFR methods against commercial FR APIs: Face++  [^1] https://www.faceplusplus.com/face-comparing/ and Luxand  [^2] https://luxand.cloud/face-api.
-(a) Red Lipstick
-(b) Blue Eyebrow
-(c) Pink Eyeshadow
-Figure 3: Comparison of concentration and hue accuracy across different methods for three makeup tasks: red lipstick, blue eyebrow, and pink eyeshadow.
-Concentration values are normalized relative to MASQUE.
-For reference-based methods, we selected images from the makeup dataset [16] used during their pre-training, that best matched the given prompt.
-For the blue eyebrow style—introduced to assess performance on an uncommon makeup—finding a suitable reference was challenging, underscoring the limitations of reference-based methods for rare styles.
-To address this, we inpainted blue eyebrows on a non-makeup image from the dataset using our framework, excluding pairwise adversarial guidance, demonstrating its capacity for targeted edits without an external reference.
+We compare MASQUE against the following baselines: TIP-IM (Yang et al., 2021), AMT-GAN (Hu et al., 2022), C2P (Shamshad et al., 2023), and DiffAM (Sun et al., 2024).
+For C2P, both its impersonating version, C2P (I), and its dodging version, C2P (D), are assessed.
+We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 (Deng et al., 2019), IRSE50 (Hu et al., 2018), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018).
+Without explicitly mentioning, we set $G=1$ in MASQUE (using a single separate guide image of the same individual).
+For reference-based methods, we selected images from the makeup dataset (Li et al., 2018) used during their pre-training that best matched the given prompt.
+In addition to public FR models, we evaluate AFR methods against two commercial APIs: Face++ [^1] https://www.faceplusplus.com/face-comparing/ and Luxand [^2] https://luxand.cloud/face-api.
 Evaluation Metric.
-We employ dodging success rate (DSR) as the primary evaluation metric for AFR.
-DSR is computed using a thresholding strategy for face verification and a closed-set strategy for face identification.
-For verification, specifically, DSR is defined as:
-$\mathrm{DSR}=\frac{1}{N}\sum_{x}\mathbb{I}\big{(}\cos(\mathrm{FR}(x_{\mathrm{p}}),\mathrm{FR}(x_{\mathrm{g}}))>\tau\big{)}\times 100\%,$
-where $\mathbb{I}$ is the indicator function, $N$ is the number of face images and $\mathrm{FR}$ represents the target face recognition model.
-$x_{\mathrm{p}}$ denotes the protected image, while $x_{\mathrm{g}}$ stands for the gallery image of the corresponding identity.
-The similarity threshold $\tau$ is set at a $0.01$ False Acceptance Rate (FAR) for each victim model.
-For face identification, we use the Rank-1 Accuracy, which measures whether the top-1 candidate list excludes the original identity of $x$.
-To assess the visual quality, we use LPIPS, PSNR, and SSIM [33].
-A lower LPIPS score reflects higher perceptual similarity, while higher PSNR and SSIM values indicate better pixel-wise and structural fidelity to the original image.
-Implementation Detail.
-MASQUE builds on the pre-trained Stable Diffusion v1.4 model, using DDIM denoising over $T=50$ steps with a fixed guidance scale of $7.5$.
-During the backward diffusion process, CA injection occurs in $[T,T-\tau_{\mathrm{attn}}]$ ($\tau_{\mathrm{attn}}=40$), localize optimization in $[T,T-\tau_{\mathrm{edit}}]$ ($\tau_{\mathrm{edit}}=5$), and adversarial guidance in $[T-\tau_{\mathrm{adv}},0]$ ($\tau_{\mathrm{adv}}=45$).
-We set $\lambda_{\mathrm{CoSi}}=0.1$, $\lambda_{\mathrm{LPIPS}}=1$, and cap optimization to $\mathrm{max}\_\mathrm{adv}=15$ iterations.
-5.2 Main Comparison Results
-Against Black-Box FR Models.
-Table 2 presents the DSR for both face verification and Rank-1 identification in black-box scenarios, evaluated using four widely used pre-trained FR feature extractors [6, 11, 25, 2].
-For each target model, the remaining three serve as surrogate models to simulate the black-box scenario, with results averaged across three makeup styles.
-Our method significantly outperforms all baselines, achieving an average DSR of $85.50\%$ for identification and $84.92\%$ for verification.
-In addition, while MASQUE without a guide image achieves a reasonable DSR, it still falls short compared to its guided counterpart.
-Furthermore, image quality degrades significantly without a guide image (see supporting evidence in Appendix B.2.).
-(a) Face++
-(b) Luxand
-Figure 4: Performance comparisons against commercial FR APIs.
-Confidence scores (lower is better) from Face++ and Luxand show that MASQUE provides stronger privacy protection than state-of-the-art noise-based and makeup-based AFR approaches.
-Real-world Performance.
-We also test our method against the commercial APIs, Face++ and Luxand, in verification mode, which assigns similarity scores from 00 to $100$.
-Figure 4 shows the results.
-As a proprietary model with unknown training data and parameters, it serves as a realistic testbed for evaluating the effectiveness of AFR methods.
-Our approach achieves the lowest similarity score across both APIs, demonstrating effectiveness in both open-source and closed-source settings, reinforcing its real-world applicability.
-TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE
-LPIPS (↓) $0.311$ $0.342$ $0.460$ $0.473$ $0.399$ $0.294$
-PSNR (↑) $32.158$ $19.512$ $18.922$ $17.990$ $18.314$ $25.815$
-SSIM (↑) $0.928$ $0.613$ $0.583$ $0.563$ $0.769$ $0.856$
-Table 3: Quantitative comparisons of the visual quality across different AFR methods. Lower values are better for LPIPS, while higher values are better for PSNR and SSIM.
+We employ DSR to evaluate the protection effectiveness of AFR methods.
+We measure visual quality using LPIPS, PSNR, and SSIM (Wang et al., 2004).
+Lower LPIPS and higher PSNR/SSIM indicate better preservation in terms of perceptual and structural similarity.
+Besides, we assess prompt following ability using metrics like concentration and hue accuracy.
+In Appendix B, we provide detailed descriptions of the hyperparameters used in MASQUE, implementation details for external image references, and formal definitions of the evaluation metrics.
+5.2 Main Results
+(a) Face++, CelebA
+(b) Luxand, CelebA
+(c) Face++, VGG-Face2
+(d) Luxand, VGG-Face2
+Figure 3: Performance comparisons between different AFR methods against two commercial APIs.
+[…]
+Table 3: Visual quality of various AFR methods on CelebA-HQ averaged across three makeup styles.
+TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE ($G=0$) MASQUE ($G=1$)
+LPIPS($\downarrow$) 0.3110.3110.311 0.3420.3420.342 0.4600.4600.460 0.4730.4730.473 0.3990.3990.399 0.2980.2980.298 0.2940.2940.294
+PSNR($\uparrow$) 32.1632.1632.16 19.5119.5119.51 18.9218.9218.92 17.9917.9917.99 18.3118.3118.31 25.3325.3325.33 25.8225.8225.82
+SSIM($\uparrow$) 0.9280.9280.928 0.6130.6130.613 0.5830.5830.583 0.5630.5630.563 0.7690.7690.769 0.8390.8390.839 0.8560.8560.856
 Visual Quality.
-Our method achieves superior image quality across multiple evaluation metrics, as summarized in Table 3 and illustrated in Figure 1.
-While TIP-IM attains the highest PSNR and SSIM due to its small perturbation budget, these pixel-level metrics often fail to reflect perceptual quality.
+MASQUE achieves superior image quality across multiple evaluation metrics, as summarized in Table 3.
+While TIP-IM attains the highest PSNR and SSIM due to its small perturbation constraint, these pixel-level metrics often fail to reflect perceptual quality (see Figure 1 for qualitative comparison results).
 In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance.
-We also confirm the advantages of MASQUE over existing AFR methods in achieving precise, localized makeup transfer (see additional experiments in Appendix B.1).
-Prompt Adherence.
-We assess prompt adherence in terms of spatial precision and color accuracy.
+Figure 4: Prompt adherence of MASQUE. Similarity score is shown at the bottom right.
+Moreover, we examine the localization capability and prompt adherence of MASQUE based on spatial precision and color accuracy.
 For spatial adherence, we use concentration, which quantifies the proportion of modifications within a binary mask of the target area.
@@
 In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility.
+In Appendix D.1, we conduct additional experiments using other distance metrics with respect to both in- and out-mask regions, illustrating the strong localization capability of MASQUE.
 These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
-(a) $\tau_{\mathrm{adv}}$
-(b) $\mathrm{max}\_\mathrm{adv}$
-Identification DSRVerification DSRLPIPS
-Figure 5: Ablation studies on hyperparameters $\tau_{\mathrm{adv}}$ and $\mathrm{max}\_\mathrm{adv}$, highlighting the trade-off between DSR and image quality.
-5.3 Ablation Studies
-Adversarial Strength vs.
-Image Quality.
-Increasing $\tau_{adv}$ enhances the adversarial effect, improving DSR but degrading image quality, as indicated by higher LPIPS values in Figure 5.
-Similarly, increasing $\mathrm{max}\_\mathrm{adv}$ amplifies protection but further compromises image quality.
-Customized Makeup. Figure 6 demonstrates the controllability of MASQUE across the following three distinct aspects. Additional visualizations are available in the Appendix C.
-Sequential Editing.
-MASQUE enables progressive makeup application, allowing users to apply makeup sequentially to multiple facial areas.
-Each additional layer of makeup enhances the protection effect, as reflected in the decreasing similarity scores shown in Figure 6.
-Intensity Control.
-The intensity of the makeup can be adjusted via text prompts.
-A simple modifier like “vivid” alters the cross-attention, resulting in color changes.
-Intensity variation has minimal effect on adversarial effectiveness.
-Style Control.
-MASQUE supports full-face transformations.
-Prompts like “tanned skin” can blend naturally, while dramatic styles like “clown makeup” are also well-adapted.
-Greater makeup coverage further strengthens protection.
-Figure 6: Examples showing the controllability of MASQUE through text prompts. The number at the bottom right indicates the similarity score between the image and the corresponding gallery image.
-6 Further Discussions
-Guide Images.
-While MASQUE with guide images enhance DSR and image quality, they may not always be available.
-As an alternative, we test self-augmented inputs as guide images via flipping and translation.
-However, preliminary results in Appendix B show limited improvements, indicating the need for more effective strategies to substitute the guide image.
-Computational Efficiency.
-MASQUE enables adversarial makeup applications without pretraining.
-However, due to the iterative nature of diffusion models, inference speed remains a computational bottleneck.
-Future work can explore acceleration schemes to improve the efficiency of MASQUE.
-Adaptive Evaluations.
-Real-world AFR threats evolve, with attackers refining strategies over time [7].
-Considering adaptive scenarios, such as post-processing transformations (e.g., compression, smoothing, and makeup removal) and adversarial fine-tuning, could help improve robustness and provide a more comprehensive evaluation of AFR methods.
+(a) Red Lipstick
+(b) Blue Eyebrow
+(c) Pink Eyeshadow
+Figure 5: Comparison of concentration and hue accuracy on CelebA-HQ across different AFR methods for three makeup styles. Concentration values are normalized relative to MASQUE.
+6 Further Analysis
+Controllable Makeup Generation.
+Figure 4 illustrates the strong controllability of MASQUE over text makeup prompts across three dimensions: sequential editing, intensity control, and style control.
+Makeup can be applied progressively to multiple facial regions, with each layer enhancing protection, as shown by decreasing similarity scores.
+Intensity is adjustable via text prompts (e.g., “vivid”), affecting appearance with minimal impact on effectiveness.
+Style prompts range from natural (e.g., “tanned skin”) to extreme (e.g., “clown makeup”), with broader coverage offering stronger protection (see Appendix E for additional visualizations of MASQUE-generated images).
+Table 4: DSR (%) of MASQUE on CelebA-HQ and “red lipstick” against various image transformations.
+Method Identification Verification
+IR152 IRSE50 MobileFace FaceNet IR152 IRSE50 MobileFace FaceNet
+MASQUE 1.001.001.00 0.970.970.97 0.940.940.94 0.710.710.71 0.970.970.97 0.950.950.95 0.690.690.69 0.660.660.66
+  + Resize 0.970.970.97 0.950.950.95 0.910.910.91 0.710.710.71 0.920.920.92 0.920.920.92 0.670.670.67 0.650.650.65
+  + Compression 1.001.001.00 0.950.950.95 0.920.920.92 0.720.720.72 0.960.960.96 0.930.930.93 0.650.650.65 0.630.630.63
+  + Gaussian Noise 0.980.980.98 0.950.950.95 0.950.950.95 0.820.820.82 0.960.960.96 0.920.920.92 0.900.900.90 0.810.810.81
+  + Blur 1.001.001.00 0.950.950.95 0.910.910.91 0.730.730.73 0.970.970.97 0.930.930.93 0.700.700.70 0.650.650.65
+  + Fog 0.950.950.95 0.970.970.97 0.930.930.93 0.880.880.88 0.920.920.92 0.970.970.97 0.840.840.84 0.810.810.81
+  + Makeup Removal 0.950.950.95 0.940.940.94 0.910.910.91 0.850.850.85 0.930.930.93 0.920.920.92 0.690.690.69 0.500.500.50
+  + DDPM Denoising 0.880.880.88 0.880.880.88 0.770.770.77 0.670.670.67 0.750.750.75 0.840.840.84 0.850.850.85 0.670.670.67
+Robustness to Image Transformation.
+We evaluate the robustness of MASQUE using the same 100100100 CelebA-HQ images and four FR models used in the main experiments, with the prompt “red lipstick”.
+We then subject the edited images to a variety of transformations, standard post-processing such as JPEG compression, blurring, natural corruptions like ImageNet-C, and adaptive attacks such as DiffAM’s makeup removal and DDPM-based denoising.
+As shown in Table 4, MASQUE maintains high dodging success rates across all cases, indicating strong resilience to real-world degradations and adaptive threats.
+This robustness may arise from the invariance of FR models, which are trained to withstand such transformations, allowing MASQUE’s protection to persist.
+Also, we test the performance of MASQUE against adversarial purification (Nie et al., 2022), where the results are presented and discussed in Appendix D.4.
+Computational Analysis.
+MASQUE does not require fine-tuning, operating directly with a pretrained Stable Diffusion model.
+In contrast, Clip2Protect fine-tunes per image (~252525s), and DiffAM performs per-dataset fine-tuning (~373737min) for each style or identity.
+At inference, MASQUE runs a one-time iterative optimization (~909090s), slightly slower than Clip2Protect (~181818s) and DiffAM (~303030s).
+Despite this, MASQUE offers key advantages: it supports arbitrary prompts and identities without retraining, making it more flexible and general.
+Under typical AFR scenarios, where protection is applied once before sharing, this adaptability outweighs the modest runtime difference.
 7 Conclusion
-We introduced MASQUE, a diffusion-based AFR framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks.
-Unlike prior AFR methods, MASQUE achieves high DSR while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse.
-Our experiments demonstrated the strong performance of MASQUE, where we highlight the important role of pairwise adversarial guidance.
-Future work may focus on real-time optimization and robustness against evolving FR models to further strengthen facial privacy protection.
+We introduced MASQUE, a diffusion-based anti-facial recognition framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks.
+Unlike prior AFR methods, MASQUE achieves high dodging success rates while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse.
+Our experiments demonstrated the superiority of MASQUE over existing baselines, where we highlight the important role of pairwise adversarial guidance.
+Future work may focus on optimizing the efficiency for real-time applications and studying the robustness against adaptively evolving FR models (Fan et al., 2025) to further strengthen MASQUE’s efficacy for facial privacy protection.
 Availability
-The code implementation of MASQUE and our experiments are open-sourced and available at https://github.com/TrustMLRG/MASQUE.
-Appendix A Related Work
-A.1 Anti-Facial Recognition
-Earlier works on AFR proposed to use obfuscation techniques to obscure the facial identity features [20] or craft $\ell_{p}$-norm bounded perturbations to fool FR models [37].
-While effective, these methods often compromise image quality, limiting their practicality in real-world scenarios.
-Poisoning-based methods [4, 27] introduced a new approach by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models.
-These techniques excel at disrupting model training or inference without visibly altering the image, but their reliance on model-specific perturbations limits their generalizability across diverse architectures.
-Unrestricted adversarial examples, which aim to create natural-looking modifications, marked a significant advancement by integrating adversarial signals into realistic alterations without relying on specific constraints.
-These strategies leverage the flexibility of generative models to balance visual coherence and adversarial effectiveness.
-Adversarial Makeup.
-Adversarial makeup [38, 12, 26, 30] has emerged as a practical solution to realize the goal of facial privacy protection.
-These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while largely preserving the aesthetic appeal.
-For example, [12] proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the generative adversarial network (GAN) framework [8].
-Recent advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect [26] and DiffAM [30], which adopt text-guided StyleGAN model and diffusion-based framework, respectively, to enable seamless adversarial face modifications with much improved visual quality.
-In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for privacy protection in dodging scenarios.
-A.2 Diffusion-Based Image Editing
-Localized Image Editing.
-Localized editing methods extend diffusion models for targeted modifications.
-For example, mask-based approaches [1, 5] constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency.
-Attention-based approaches [9, 32] guide edits via attention injection, achieving better global structure preservation but suffering from unintended changes (editing leakage).
-Our proposed method, MASQUE, utilizes both mask-based and attention-based guidance without any fine-tuning, combining mask precision with attention flexibility, ensuring semantically consistent and region-specific modifications while addressing the limitations of prior approaches.
-Text-Guided Diffusion Models.
-Text-guided diffusion models are a major advancement in generative AI, capable of synthesizing high-quality images from natural language descriptions.
-Built on denoising diffusion probabilistic models (DDPMs) [10], they iteratively refine random noise into coherent images guided by text prompts.
-Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation.
-Mask-based approaches [1, avrahami2022blended, 5] achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions.
-While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios.
-On the other hand, mask-free attention-based methods [9, cao2023masactrl, 32] use attention injection mechanisms to guide edits without requiring explicit masks.
-These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region.
-In this work, we leverage mask-based cross-attention guidance [18] to achieve localized adversarial perturbations in the form of makeup.
-This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts while addressing the limitations of structural inconsistencies seen in previous methods.
-A.3 Makeup Transfer
-Advancements in makeup transfer have moved from GAN-based methods to more advanced diffusion models.
-Traditional approaches like BeautyGAN [16] and PSGAN [17] use histogram matching, attention mechanisms, and spatial encodings to transfer makeup while preserving structure.
-CPM [21] expands beyond color transfer using UV mapping, while RamGAN [35] ensure component consistency.
-However, GAN-based models struggle with extreme styles and rely on imprecise pseudo-paired data, limiting fidelity.
-To overcome these challenges, diffusion-based methods like SHMT [31] and Stable-Makeup [40] offer improved realism and robustness.
-Specifically, SHMT eliminates reliance on pseudo-paired data through a self-supervised “decoupling-and-reconstruction” framework, leveraging a Laplacian pyramid for hierarchical texture transfer.
-Stable-Makeup employs a diffusion model with a Detail-Preserving encoder and cross-attention layers to ensure precise, structure-preserving makeup application.
-These approaches set a new standard for high-fidelity transfer.
-However, our study takes a different approach—rather than transferring makeup from a reference image, we focus on text-guided makeup generation using diffusion models, enabling intuitive and flexible image editing to create customizable makeup based solely on textual descriptions.
-This also eliminates the need for model fine-tuning on specific makeup images, allowing us to generate the desired makeup directly during inference, resulting in a more efficient and adaptable approach.
-Appendix B Additional Experiments
-B.1 Localized Editing
-To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask, isolating the corresponding areas in both images via element-wise multiplication.
-The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons.
-Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences.
-Metric Method In-Mask Out-Mask $\textbf{Diff}\ \bm{\Delta}$
-DISTS ()($\downarrow$)( ↓ ) TIP-IM $2.013$ $0.094$ $1.919$
-AMT-GAN $7.993$ $0.149$ $7.844$
-C2P (I) $8.945$ $0.153$ $8.792$
-C2P (D) $10.011$ $0.169$ $9.842$
-DiffAM $9.645$ $0.141$ $9.504$
-MASQUE $12.306$ $0.104$ $12.202$
-LPIPS ()($\downarrow$)( ↓ ) TIP-IM $0.322$ $0.305$ $0.017$
-AMT-GAN $0.360$ $0.316$ $0.044$
-C2P (I) $0.400$ $0.447$ $-0.047$
-C2P (D) $0.436$ $0.460$ $-0.025$
-DiffAM $0.509$ $0.410$ $0.099$
-MASQUE $0.434$ $0.290$ $0.144$
-PieAPP ()($\downarrow$)( ↓ ) TIP-IM $3.621$ $0.462$ $3.1591$
-AMT-GAN $21.325$ $0.814$ $20.511$
-C2P (I) $17.186$ $1.138$ $16.049$
-C2P (D) $20.850$ $1.314$ $19.536$
-DiffAM $17.523$ $1.312$ $16.211$
-MASQUE $21.433$ $0.645$ $20.789$
-Table 4: Comparison of perceptual similarity metrics across in-mask and out-mask regions and different AFR methods.
-Table 4 compares the effectiveness of localized edits, where higher $\Delta$ values indicate stronger localization of perturbations.
-TIP-IM [37], as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions.
-In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
-B.2 Pairwise Adversarial Guidance
-MASQUE introduces pairwise adversarial guidance to protect identities without external target references, distinguishing it from impersonation-oriented methods.
-By aligning features between the guide and original images, we identify identity-relevant cues and inject subtle adversarial signals that steer the diffusion process away from a recognizable identity manifold.
-The guide image serves as a proxy for the gallery image, providing guidance on the direction in which the latent representation should be altered.
-This eliminates reliance on external identities, mitigating ethical concerns related to impersonation while ensuring identity protection remains non-intrusive.
-Leveraging multiple guide images further stabilizes adversarial training, reducing bias toward any single representation and improving robustness.
-To evaluate its effectiveness, we use FaceNet as the target model and select $100$ identities from CelebA-HQ, varying the number of guide images per identity.
-These evaluation identities are distinct from the main experiments, ensuring at least $12$ images per identity to accommodate setups with up to $10$ guide images.
-Effectiveness is measured using the confidence score, the gap between the similarity score of the original identity’s gallery image and that of the misidentified identity.
-Table 5 shows that increasing guide images enhances identity obfuscation, leading to higher confidence scores while reducing similarity with the original identity.
-Moreover, images generated with guide images in the pairwise adversarial guidance step exhibit higher visual quality than those relying solely on adversarial loss on the original image.
-This confirms that simply maximizing distance from the original image is ineffective, as it disrupts the diffusion process’s goal of preserving natural image characteristics, leading to instability in both image quality and adversarial performance.
-These results demonstrate that MASQUE’s pairwise adversarial guidance provides a robust, privacy-centric solution, balancing strong identity obfuscation with high visual fidelity.
-#Guide Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) Confidence()($\uparrow$)( ↑ ) Similarity()($\downarrow$)( ↓ ) LPIPS()($\downarrow$)( ↓ )
-00 $0.66$ $0.60$ $0.114$ $0.355$ $0.303$
-$1$ $0.67$ $0.63$ $0.135$ $0.349$ $0.292$
-$2$ $0.72$ $0.71$ $0.165$ $0.297$ $0.291$
-$5$ $0.74$ $0.74$ $0.177$ $0.283$ $0.291$
-$10$ $0.76$ $0.77$ $0.197$ $0.266$ $0.291$
-Table 5: Comparison of identification confidence, verification similarity, and image quality metrics across different numbers of guide images employed in MASQUE.
-Guide Method Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ )
-Pairwise ($G=1$) $0.72$ $0.66$ $25.691$
-Self-Guidance ($G=0$) $0.56$ $0.40$ $25.812$
-    + Horizontal Flip $0.57$ $0.50$ $25.834$
-    + Vertical Flip $0.12$ $0.06$ $26.566$
-    + Translation $0.55$ $0.38$ $25.807$
-Table 6: Impact of adversarial image guidance in DSR and PSNR. We test pairwise- and self-guidance with different augmentations.
-B.3 Self-Augmented Guidance
-To address cases where users may lack suitable guide images, we tested using self-augmented input images as guide images.
-This experiment was conducted on $100$ identities from CelebA-HQ with the prompt “red lipstick”, targeting the FaceNet model.
-For augmentation, we applied horizontal flipping, vertical flipping, and random translation with a maximum shift of $20$ pixels.
-Results, as shown in Table 6, compare these approaches against the input image itself as the guide image.
-Horizontal flipping and random translation perform similarly to using the input image as the guide, whereas vertical flipping shows minimal adversarial effect.
-This is likely because vertically flipped faces are hard to align with facial images by the recognition model.
-B.4 DSR under Rank-5 Accuracy
-We also report DSR for the face identification task, using rank-5 accuracy to assess dodging success.
-The results are presented in Table 7.
-Notably, MASQUE achieves the highest DSR in both rank-1 and rank-5 settings, demonstrating its effectiveness in misleading identification models.
-AMT-GAN DiffAM C2P (I) C2P (D) MASQUE
-Rank-1 $54.39$ $61.25$ $36.92$ $71.17$ $92.08$
-Rank-5 $28.58$ $45.58$ $18.25$ $48.67$ $81.42$
-Table 7: DSR with Rank-1 and Rank-5 metrics across AFR methods
-B.5 Mask and Perturbation Localization
-Figure 7: Visualizations of MASQUE with and without masking. The first row uses prompt “shimmery gold eyeshadow”, while the second row uses “blue eyeshadow”.
-To understand the impact of masking in MASQUE, we conduct an ablation study with and without applying a mask.
-Figure 7 shows that the absence of a mask causes the model to struggle with spatial control, leading to unconstrained perturbations across the entire image rather than confining makeup to the intended area.
-Appendix C Additional Visualizations
-Figure 8: Visualizations of MASQUE-generated images with or without using the edit loss.
-Additional visual results of image-guided makeup transfer are provided in Figure 8 to accommodate space limitations in the main text.
-MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics.
-To evaluate color variation, we provide examples featuring both a male and a female subject.
-Additionally, by testing prompts that target different facial regions (eyes, lips, cheeks, skin, and full-face makeup), we demonstrate MASQUE’s ability to generate realistic makeup across various regions.
-We further provide visualizations without $\mathcal{L}_{\text{edit}}$ optimization to illustrate its impact.
-Appendix D Algorithm Pseudocode
-Input: $\bm{z}_{\mathrm{edit}},\bm{z}_{\mathrm{rec}}$: Edited and original latent, $\bm{e}_{\mathrm{edit}},\bm{e}_{\mathrm{rec}}$: Edited and original text embeddings, $\mathcal{D}$: Stable Diffusion model, $T$: Number of diffusion steps, $\tau_{\mathrm{attn}},\tau_{\mathrm{edit}},\tau_{\mathrm{adv}}$: Step thresholds, $\mathcal{M}$: Binary region mask, $\mathcal{I}_{\mathrm{target}}$: Target token indices, $d_{k}$: Key/Query dimensionality, $\lambda_{\mathrm{edit}},\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$: Loss weights, $\eta$: Learning rate, $\sigma(t)$: Noise scale function, $\bm{x}$: Input image, $\tilde{\bm{x}}$: Guide image, $max\_adv$: Maximum adversarial iterations
+The code implementations of our method and for reproducing our experimental results are open-sourced and available at: https://github.com/TrustMLRG/MASQUE.
+Appendix A Algorithm Pseudocode
+Algorithm 1 presents the pseudocode of MASQUE, which consists of modules such as precise null-text inversion, cross-attention fusion with masking, and pairwise adversarial guidance.
+Input: $\bm{z}_{\mathrm{edit}},\bm{z}_{\mathrm{rec}}$: edited and original latent, $\bm{e}_{\mathrm{edit}},\bm{e}_{\mathrm{rec}}$: edited and original text embeddings, $\mathcal{D}$: Stable Diffusion model, $T$: number of diffusion steps, $\tau_{\mathrm{attn}},\tau_{\mathrm{edit}},\tau_{\mathrm{adv}}$: step thresholds, $\mathcal{M}$: binary region mask, $\mathcal{I}_{\mathrm{target}}$: target token indices, $d_{k}$: key/query dimensionality, $\lambda_{\mathrm{edit}},\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$: loss weights, $\eta$: learning rate, $\sigma(t)$: noise scale function, $\bm{x}$: input image, $\tilde{\bm{x}}$: guide image, $m_{\mathrm{adv}}$: maximum adversarial iterations
 Output: Final protected image $\bm{x}_{\mathrm{p}}$
-for $k\leftarrow T$ to $1$ do
+for $k\leftarrow T$ to 111 do
  if $k>T-\tau_{\mathrm{attn}}$ then
@@
     ${}\mathrm{CA}^{\mathrm{refined}}\leftarrow\{\}$
-            for $\ell\in\text{CrossAttnLayers}$ do
-                   $Q_{\ell},K_{\ell},V_{\ell}\leftarrow\mathcal{D}.\mathrm{UNet.GetCrossAttentionComponents}(\bm{z}_{\mathrm{edit}},\bm{e}_{\mathrm{edit}},\ell)$
-                  $\mathrm{CA}_{\ell}\leftarrow\mathrm{Softmax}\Bigl{(}\frac{Q_{\ell}\,K_{\ell}^{\top}}{\sqrt{d_{k}}}\Bigr{)}$
-                  $\mathrm{CA}_{\ell}^{\mathrm{refined}}\leftarrow\mathrm{CA}_{\ell}\cdot V_{\ell}$
-                  $\mathrm{CA}^{\mathrm{refined}}\leftarrow\mathrm{CA}^{\mathrm{refined}}\cup\{\mathrm{CA}_{\ell}^{\mathrm{refined}}\}$
+    for $l\in\text{CrossAttnLayers}$ do
+       $Q_{l},K_{l},V_{l}\leftarrow\mathcal{D}.\mathrm{UNet.GetCrossAttentionComponents}(\bm{z}_{\mathrm{edit}},\bm{e}_{\mathrm{edit}},l)$
+       $\mathrm{CA}_{l}\leftarrow\mathrm{Softmax}\Bigl{(}\frac{Q_{l}\,K_{l}^{\top}}{\sqrt{d_{k}}}\Bigr{)}$
+       $\mathrm{CA}_{l}^{\mathrm{refined}}\leftarrow\mathrm{CA}_{l}\cdot V_{l}$
+       $\mathrm{CA}^{\mathrm{refined}}\leftarrow\mathrm{CA}^{\mathrm{refined}}\cup\{\mathrm{CA}_{l}^{\mathrm{refined}}\}$
  if $k>T-\tau_{\mathrm{edit}}$ then
@@
     $\mathcal{L}_{\mathrm{edit}}\leftarrow 0$
-            foreach $\mathrm{CA}_{\ell}^{\mathrm{refined}}\in\mathrm{CA}^{\mathrm{refined}}$ do
-                   $\mathrm{CA}_{\mathrm{target}}\leftarrow\mathrm{ExtractTargetAttention}\Bigl{(}\mathrm{CA}_{\ell}^{\mathrm{refined}},\mathcal{I}_{\mathrm{target}}\Bigr{)}$
+    foreach $\mathrm{CA}_{l}^{\mathrm{refined}}\in\mathrm{CA}^{\mathrm{refined}}$ do
+       $\mathrm{CA}_{\mathrm{target}}\leftarrow\mathrm{ExtractTargetAttention}\Bigl{(}\mathrm{CA}_{l}^{\mathrm{refined}},\mathcal{I}_{\mathrm{target}}\Bigr{)}$
        $\mathrm{CA}_{\mathrm{masked}}\leftarrow\mathrm{CA}_{\mathrm{target}}\odot\mathcal{M}$
@@
     //Perform pairwise adversarial optimization
-            for $i\leftarrow 0$ to $max\_adv-1$ do
+    for $i\leftarrow 0$ to $m_{\mathrm{adv}}-1$ do
        $\bm{x}_{\mathrm{rec}}\leftarrow\mathcal{D}.\mathrm{VAE}.\mathrm{Decode}(\bm{z}_{\mathrm{edit}})$
@@
 return $\bm{x}_{\mathrm{p}}$
-Algorithm 1 Adversarial Makeup Generation with CA Guidance
+Algorithm 1 Adversarial Makeup Generation with Pairwise Adversarial CA Guidance
+Appendix B Detailed Experimental Settings
+Implementation Details.
+MASQUE builds on the pre-trained Stable Diffusion v1.4 model (Rombach et al., 2022), using a denoising diffusion implicit model (DDIM) (Song et al., 2020) denoising over $T=50$ steps with a fixed guidance scale of 7.57.57.5.
+During backward diffusion, CA injection occurs in $[T,T-\tau_{\mathrm{attn}}]$ ($\tau_{\mathrm{attn}}=40$), localize optimization in $[T,T-\tau_{\mathrm{edit}}]$ ($\tau_{\mathrm{edit}}=5$), and adversarial guidance in $[T-\tau_{\mathrm{adv}},0]$ ($\tau_{\mathrm{adv}}=45$).
+We set $\lambda_{\mathrm{CoSi}}=0.1$, $\lambda_{\mathrm{LPIPS}}=1$, and cap optimization to $m_{\mathrm{adv}}=15$ iterations.
+We select the value of these hyperparameters based on the prior literature or our ablation studies presented in Appendix C.3.
+Reference Images.
+For reference-based methods, we selected images from the makeup dataset (Li et al., 2018) used during their pre-training that best matched the given prompt.
+For the “blue eyebrow” style—introduced to assess performance on an uncommon makeup—finding a suitable reference was challenging, underscoring the limitations of reference-based methods for rare styles.
+To address this, we inpainted blue eyebrows on a non-makeup image from the dataset using our framework, excluding pairwise adversarial guidance, showing its capacity for targeted edits without an external reference.
+Evaluation Metrics.
+We use the dodging success rate (DSR) as the primary metric for evaluating the protection effectiveness of AFR.
+DSR is computed using a thresholding strategy for face verification and a closed-set strategy for face identification.
+For verification, DSR is defined as:
+$\mathrm{DSR}_{\mathrm{FV}}=\frac{1}{|\mathcal{D}|}\sum_{(\bm{x},y)\in\mathcal{D}}\mathds{1}\big{(}\cos(f(\bm{x}_{\mathrm{p}}),f(\bm{x}_{\mathrm{g}}))>\gamma\big{)},$
+where $f$ denotes the feature extractor of the FR model, and $\bm{x}_{\mathrm{g}}$ stands for the gallery image of the victim identity.
+The similarity threshold $\gamma$ is set at a 0.010.010.01 false acceptance rate (FAR) for each FR model.
+For face identification, the DSR of an AFR method is defined as the optimization objective of Equation 1.
+In particular, we use the rank-1 accuracy, which measures whether the top-1 candidate list excludes the original identity corresponding to the victim user’s image $\bm{x}$.
+Table 5: Comparison of identification confidence, verification similarity, and image quality metrics across different numbers of guide images employed in MASQUE.
+#Guide Images Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) Confidence()($\uparrow$)( ↑ ) Similarity()($\downarrow$)( ↓ ) LPIPS()($\downarrow$)( ↓ )
+0 0.660.660.66 0.600.600.60 0.1140.1140.114 0.3550.3550.355 0.3030.3030.303
+111 0.670.670.67 0.630.630.63 0.1350.1350.135 0.3490.3490.349 0.2920.2920.292
+222 0.720.720.72 0.710.710.71 0.1650.1650.165 0.2970.2970.297 0.2910.2910.291
+555 0.740.740.74 0.740.740.74 0.1770.1770.177 0.2830.2830.283 0.2910.2910.291
+101010 0.760.760.76 0.770.770.77 0.1970.1970.197 0.2660.2660.266 0.2910.2910.291
+Appendix C Ablation Studies
+In this section, we present detailed ablation studies to comprehensively examine the impact of different components of our approach.
+All experiments are conducted on the CelebA-HQ dataset, using the FaceNet model as the target and “red lipstick” as the makeup prompt.
+C.1 Number of Images for Adversarial Guidance
+MASQUE introduces pairwise adversarial guidance to protect identities without external target references, distinguishing it from impersonation-oriented methods.
+By aligning features between the guide and original images, we identify identity-relevant cues and inject subtle adversarial signals that steer the diffusion process away from a recognizable identity manifold.
+The guide image serves as a proxy for the gallery image, providing guidance on the direction in which the latent representation should be altered.
+This eliminates reliance on external identities, mitigating ethical concerns related to impersonation while ensuring identity protection remains non-intrusive.
+Using multiple guide images further stabilizes training, reducing bias toward any single representation and improving robustness.
+To evaluate its effectiveness, we use FaceNet as the target model and select 100100100 identities from CelebA-HQ, varying the number of guide images per identity.
+These evaluation identities are distinct from the main experiments, ensuring at least 121212 images per identity to accommodate setups with up to 101010 guide images.
+Effectiveness is measured using the confidence score, the gap between the similarity score of the original identity’s gallery image and that of the misidentified identity.
+Table 5 shows that increasing guide images enhances identity obfuscation, leading to higher confidence scores while reducing similarity with the original identity.
+Moreover, images generated with guide images in the pairwise adversarial guidance step exhibit higher visual quality than those relying solely on adversarial loss on the original image.
+This confirms that simply maximizing distance from the original image is ineffective, as it disrupts the diffusion process’s goal of preserving natural image characteristics, leading to instability in both image quality and adversarial performance.
+These results demonstrate that MASQUE’s pairwise adversarial guidance provides a robust, privacy-centric solution, balancing strong identity obfuscation with high visual fidelity.
+[…]
+Method #Guide Images Augmentation Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ )
+Baseline $G=1$ none 0.720.720.72 0.660.660.66 25.6925.6925.69
+Baseline $G=0$ none 0.560.560.56 0.400.400.40 25.8125.8125.81
+Self-aug. $G=0$ horizontal flip 0.570.570.57 0.500.500.50 25.8325.8325.83
+Self-aug. $G=0$ vertical flip 0.120.120.12 0.060.060.06 26.5726.5726.57
+Self-aug. $G=0$ translation 0.550.550.55 0.380.380.38 25.8125.8125.81
+(a) Step Threshold $\tau_{\mathrm{adv}}$
+(b) Opt. Steps $m_{\mathrm{adv}}$
+(c) CoSi Weight $\lambda_{\mathrm{CoSi}}$
+(d) LPIPS Weight $\lambda_{\mathrm{LPIPS}}$
+Figure 6: Ablations on MASQUE showing trade-off between protection strength and image quality.
+C.2 Self-Augmented Guidance
+To address cases where users may lack suitable guide images, we tested using self-augmented input images.
+For augmentation, we applied horizontal flipping, vertical flipping, and random translation with a maximum shift of 202020 pixels.
+Results, as shown in Table 6, compare these approaches against the input image itself as the guide image.
+Horizontal flipping and random translation perform similarly to using the input image as the guide, whereas vertical flipping shows minimal adversarial effect.
+This is likely because vertically flipped faces are hard to align with facial images for the recognition model.
+C.3 Adversarial Strength versus Image Quality
+Figure 6 presents our ablation studies on the impact of hyperparameters used in Algorithm 1: $\tau_{\mathrm{adv}}$, $m_{\mathrm{adv}}$, $\lambda_{\mathrm{CoSI}}$, and $\lambda_{\mathrm{LPIPS}}$.
+Increasing the value of $\tau_{\mathrm{adv}}$ enhances the adversarial effect, improving DSR but degrading image quality, as indicated by higher LPIPS values.
+Similarly, increasing the number of pairwise adversarial optimization steps $m_{\mathrm{adv}}$ amplifies protection but further compromises image quality.
+According to Figures 6(c) and 6(d), we select $\lambda_{\mathrm{CoSi}}=0.1$, $\lambda_{\mathrm{LPIPS}}=1.0$ to balance the trade-off between the dodging success rates and the visual quality of generated images.
+Figure 7: Visualizations of MASQUE ($G=1$) with and without masking. The first row uses prompt “shimmery gold eyeshadow”, while the second row uses “blue eyeshadow”.
+C.4 Mask and Perturbation Localization
+To understand the impact of masking, we further conduct an ablation study with and without applying a mask in the proposed framework of MASQUE.
+Figure 7 shows that the absence of a mask causes the model to struggle with spatial control, leading to unconstrained perturbations across the entire image rather than confining makeup to the intended area.
+Table 7: Comparison of similarity metrics between AFR methods across in- and out-mask regions.
+Metric Type AFR Method
+TIP-IM AMT-GAN C2P (I) C2P (D) DiffAM MASQUE ($G=1$)
+DISTS ()($\downarrow$)( ↓ ) In-Mask 2.0132.0132.013 7.9937.9937.993 8.9458.9458.945 10.01110.01110.011 9.6459.6459.645 12.30612.30612.306
+Out-Mask 0.0940.0940.094 0.1490.1490.149 0.1530.1530.153 0.1690.1690.169 0.1410.1410.141 0.1040.1040.104
+Diff ($\Delta$) 1.9191.9191.919 7.8447.8447.844 8.7928.7928.792 9.8429.8429.842 9.5049.5049.504 12.20212.20212.202
+LPIPS ()($\downarrow$)( ↓ ) In-Mask 0.3220.3220.322 0.3600.3600.360 0.4000.4000.400 0.43560.43560.4356 0.5090.5090.509 0.4340.4340.434
+Out-Mask 0.3050.3050.305 0.3160.3160.316 0.4470.4470.447 0.4600.4600.460 0.4100.4100.410 0.2900.2900.290
+Diff ($\Delta$) 0.0170.0170.017 0.0440.0440.044 0.047-0.047- 0.047 0.025-0.025- 0.025 0.0990.0990.099 0.1440.1440.144
+PieAPP ()($\downarrow$)( ↓ ) In-Mask 3.6213.6213.621 21.32521.32521.325 17.18617.18617.186 20.85020.85020.850 17.52317.52317.523 21.43421.43421.434
+Out-Mask 0.4620.4620.462 0.8130.8130.813 1.1381.1381.138 1.3141.3141.314 1.3121.3121.312 0.6450.6450.645
+Diff ($\Delta$) 3.1593.1593.159 20.51120.51120.511 16.04916.04916.049 19.53619.53619.536 16.21116.21116.211 20.78920.78920.789
+[…]
+Method AMT-GAN DiffAM C2P(I) C2P(D) MASQUE ($G=1$)
+Rank-1 Accuracy 54.3954.3954.39 61.2561.2561.25 36.9236.9236.92 71.1771.1771.17 92.0892.0892.08
+Rank-5 Accuracy 28.5828.5828.58 45.5845.5845.58 18.2518.2518.25 48.6748.6748.67 81.4281.4281.42
+Table 9: Visual assessment of AFR methods on VGG-Face2-HQ averaged for three makeup styles.
+TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE ($G=1$)
+LPIPS($\downarrow$) 0.2200.2200.220 0.2750.2750.275 0.3570.3570.357 0.3560.3560.356 0.4990.4990.499 0.2050.2050.205
+PSNR($\uparrow$) 26.5826.5826.58 20.1820.1820.18 19.1219.1219.12 19.3719.3719.37 14.2914.2914.29 25.3425.3425.34
+SSIM($\uparrow$) 0.8820.8820.882 0.7970.7970.797 0.6640.6640.664 0.6730.6730.673 0.4510.4510.451 0.8720.8720.872
+[…]
+Dataset Identification Verification
+IR152 IRSE50 MobileFace FaceNet IR152 IRSE50 MobileFace FaceNet
+CelebA-HQ 0.280.280.28 0.360.360.36 0.550.550.55 0.280.280.28 0.120.120.12 0.100.100.10 0.180.180.18 0.170.170.17
+VGGFace2-HQ 0.880.880.88 0.800.800.80 0.830.830.83 0.770.770.77 0.860.860.86 0.820.820.82 0.920.920.92 0.780.780.78
+Appendix D Additional Experiments and Discussions
+D.1 Localized Editing
+To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask.
+The corresponding areas in both images are isolated via element-wise multiplication.
+The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons.
+Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences.
+Table 7 compares the effectiveness of localized edits, where higher $\Delta$ values indicate stronger localization of perturbations.
+TIP-IM (Yang et al., 2021), as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions.
+In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
+D.2 DSR under Rank-5 Accuracy
+We also report DSR for the face identification task, using rank-5 accuracy to assess dodging success.
+The results are presented in Table 8.
+Notably, MASQUE achieves the highest DSR in both rank-1 and rank-5 settings, demonstrating its effectiveness in misleading identification models.
+D.3 VGG-Face2 Image Quality
+We report image quality metrics on the VGG-Face2 dataset, averaged over four face recognition models and three makeup styles.
+The results are shown in Table 9.
+MASQUE demonstrates consistently strong performance, in line with the results observed on the CelebA-HQ dataset.
+D.4 Robustness against Adversarial Purification
+Table 10 assesses MASQUE against adversarial purification techniques (Nie et al., 2022).
+When the purifier is fine-tuned on CelebA-HQ, it successfully removes MASQUE-generated perturbations on images from the same dataset.
+However, its effectiveness drops significantly on images from VGGFace2-HQ, revealing limited generalization to unseen identities and domains.
+Moreover, the computational overhead of adversarial purification—especially for high-resolution inputs—makes it impractical for real-time face recognition systems, particularly in tracking or surveillance scenarios.
+D.5 Limitations and Broader Impact
+MASQUE may fail when the mask is very small, as downscaling may shrink or remove it entirely.
+We applied iterative dilation during early diffusion steps to address this, but very small masks may still vanish.
+Moreover, real-world AFR threats are constantly evolving, with adversaries continuously refining their strategies (Fan et al., 2025).
+Incorporating adaptive threat models and adversarial fine-tuning into the evaluation process could improve robustness and lead to a more comprehensive assessment of defense performance.
+While our method is designed to defend against unauthorized or invasive FR systems, it may raise dual-use concerns.
+Specifically, the same techniques could potentially be misused to evade legitimate or authorized FR systems, such as those used in security or law enforcement contexts.
+We believe it is important to anticipate and acknowledge these risks, and we encourage future work to consider safeguards or contextual constraints that limit misuse while preserving individual privacy protections.
+Appendix E Additional Visualizations
+Additional visual results of image-guided makeup transfer are provided in Figure 8 and Figure 9 to accommodate space limitations in the main text.
+MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics.
+Figure 8: Visualizations of MASQUE-generated adversarial images from the VGG-Face2-HQ Dataset.
+To evaluate color variation, we provide examples featuring both a male and a female subject.
+Additionally, by testing prompts that target different facial regions (eyes, lips, cheeks, skin, and full-face makeup), we demonstrate MASQUE’s ability to generate realistic makeup across various regions.
+We further provide visualizations without $\mathcal{L}_{\text{edit}}$ optimization to illustrate its impact.
+Figure 9: Visualizations of MASQUE-generated adversarial images with or without using the edit loss.